### v1

# WorldForge: Unlocking Emergent 3D/4D Generation in Video Diffusion Model via Training-Free Guidance

Abstract
Recent video diffusion models demonstrate strong potential in spatial intelligence tasks due to their rich latent world priors. However, this potential is hindered by their limited controllability and geometric inconsistency, creating a gap between their strong priors and their practical use in 3D/4D tasks. As a result, current approaches often rely on retraining or fine-tuning, which risks degrading pretrained knowledge and incurs high computational costs. To address this, we propose WorldForge, a training-free, inference-time framework composed of three tightly coupled modules. Intra-Step Recursive Refinement introduces a recursive refinement mechanism during inference, which repeatedly optimizes network predictions within each denoising step to enable precise trajectory injection. Flow-Gated Latent Fusion leverages optical flow similarity to decouple motion from appearance in the latent space and selectively inject trajectory guidance into motion-related channels. Dual-Path Self-Corrective Guidance compares guided and unguided denoising paths to adaptively correct trajectory drift caused by noisy or misaligned structural signals. Together, these components inject fine-grained, trajectory-aligned guidance without training, achieving both accurate motion control and photorealistic content generation. Extensive experiments across diverse benchmarks validate our method’s superiority in realism, trajectory consistency, and visual fidelity. This work introduces a novel plug-and-play paradigm for controllable video synthesis, offering a new perspective on leveraging generative priors for spatial intelligence.
Figure 1: We present WorldForge, a fully training-free framework leveraging a pre-trained video diffusion model for various 3D/4D tasks, such as monocular 3D scene generation (up) and dynamic 4D scene re-rendering (down), enabling precise camera trajectory control and high-quality outputs.
1 Introduction
Recent advances in generative modeling, particularly video diffusion models (VDM) (Blattmann et al., 2023; Wan et al., 2025; Yang et al., 2024; Google DeepMind, 2025), have greatly expanded the scope and capabilities of spatial intelligence (Cao et al., 2025) tasks such as 3D and 4D scene understanding (Bahmani et al., 2025a; b), reconstruction (Wang et al., 2025a; Wu et al., 2025; Shi et al., 2024), and generation (Yu et al., 2024c; 2025). Trained on massive, diverse video datasets, these models inherently encode rich spatiotemporal priors that capture structural, temporal, and motion-related patterns. Leveraging such priors offers significant advantages for achieving realistic and coherent spatial transformations, enabling applications in novel view synthesis (You et al., 2025; Xiao et al., 2025), panoramic video generation (Wang et al., 2024b; Ma et al., 2024a), 3D scene generation (Liu et al., 2024; Liang et al., 2025; Wang et al., 2024c), and dynamic scene reconstruction (Bai et al., 2025a; Yu et al., 2025; Van Hoorick et al., 2024). Moreover, VDMs are increasingly used to construct “world models” (Bar et al., 2025; Duan et al., 2025; Bruce et al., 2024), understood as structured internal representations of physical environments that support predictive reasoning, planning, and decision-making in embodied AI systems.
Despite their strong generative capabilities, current video diffusion models face fundamental limitations, including limited controllability, spatial–temporal consistency, and geometric fidelity, particularly when applied to 3D or 4D tasks (Wang et al., 2024c; He et al., 2024; Ling et al., 2024; Xing et al., 2024). While they can be loosely guided with text prompts or reference frames (Kong et al., 2024; Wan et al., 2025), they generally lack the ability to follow precise user-defined motion constraints, such as a specified 6-DoF camera trajectory or object pose evolution (Hu, 2024; Ma et al., 2024b). This lack of trajectory-level control is especially problematic in tasks requiring spatial consistency across views, such as novel view synthesis or free-viewpoint rendering. Furthermore, these models often entangle scene and camera motion, making it difficult to separate object dynamics from viewpoint changes (Yu et al., 2024c; Liu et al., 2024). Consequently, attempts to synthesize videos with fixed scenes or smooth camera paths often result in unintended object deformations or scene instability. These limitations hinder their applicability in domains requiring structured spatial reasoning or controllable video generation.
To handle these limitations, prior works (Jeong et al., 2025; Ren et al., 2025; Yu et al., 2025; Zhang et al., 2025) have explored two main directions. The first involves training or fine-tuning the generative backbone on multiview or motion-conditioned data, often with explicit modules to encode trajectory information (Bai et al., 2025a; Xiao et al., 2024; Bai et al., 2025b). While this approach can improve alignment and control, it is computationally costly, may generalize poorly to diverse scenes, and risks degrading the model’s pretrained priors during fine-tuning. The second line of work adopts a “warping-and-repainting” strategy (Ma et al., 2025b; Liu et al., 2025; Ma et al., 2025a; You et al., 2025), in which input frames are lifted into a partial 3D representation (e.g., via depth estimation (Piccinelli et al., 2024; Yin et al., 2023)), re-projected along a user-defined camera path, and then refined by a generative model to fill missing regions. Although more flexible, such methods face notable robustness limitations, as pretrained models are not designed to process out-of-distribution (OOD) (Yu et al., 2024a) inputs such as warped or disoccluded images. As a result, they often produce artifacts and missing structures (e.g., suspended components or fragmented geometry). Moreover, their bias toward dynamic training data leads to hallucinated motion even in static scenes, undermining view consistency. In summary, the combined effects of OOD inputs and dynamic-data bias make it challenging to balance fine-grained controllability with generation quality and generalization, which remains an open problem.
To address this challenge, we aim to inject precise control into VDMs while preserving their valuable priors. For this purpose, we propose a general inference-time guidance paradigm that leverages the rich priors of large-scale VDMs (Blattmann et al., 2023; Wan et al., 2025) in spatial intelligence tasks, such as geometry-aware 3D scene generation and video trajectory control. Our method adopts a warping-and-repainting pipeline, in which input frames are warped along a reference trajectory and then used as conditional inputs in the repainting stage. Building on this, we develop a unified, training-free framework composed of three complementary mechanisms, each designed to address a specific challenge in trajectory-controlled generation.
First, to ensure the generated motion follows the target trajectory derived from depth-based rendering (Wang et al., 2025b; Piccinelli et al., 2024), we introduce Intra-Step Recursive Refinement (IRR). It embeds a micro-scale predict–correct loop within each denoising step: before the next timestep, predicted content in observed regions is replaced with the corresponding ground-truth observations. This incremental correction allows trajectory control signals to be injected at every step, enabling fine-grained, stepwise guidance that keeps the motion aligned with the target trajectory.
Second, we observe that different channels of the VAE-encoded (Kingma & Welling, 2013; Foti et al., 2022) latent representation encode different information, with some channels specializing in appearance and others in motion. Directly overwriting all channels when injecting trajectory signals can inadvertently degrade fine-grained visual details. To address this, we propose Flow-Gated Latent Fusion (FLF), which leverages optical-flow similarity to selectively and dynamically inject trajectory information only into motion-relevant channels, while leaving appearance-relevant channels unmodified. This selective modulation effectively decouples appearance from motion, allowing for precise viewpoint manipulation while preserving content fidelity.
Finally, while warping-based rendering effectively enforces user-defined trajectories in view synthesis, it inevitably introduces noise stemming from imperfect depth estimation, occlusions, and scene misalignments. These imperfections often lead to visual artifacts such as ghosting, structural distortions, and degraded temporal coherence (You et al., 2025). To better balance trajectory adherence and generation fidelity, we propose a Dual-Path Self-Corrective Guidance (DSG) strategy. Inspired by CFG (Ho & Salimans, 2021), DSG introduces two concurrent denoising pathways during inference: a non-guided path that relies on the model’s internal priors and yields high-fidelity but uncontrolled outputs, and a guided path conditioned on the warped input trajectory, which enforces camera motion but is prone to artifacts. By utilizing the difference between these two paths at each denoising step, DSG computes a dynamic correction term that softly adjusts the guided path toward the perceptual quality of the non-guided path. This self-corrective mechanism effectively mitigates trajectory-induced degradation while maintaining alignment with the target camera path, thereby improving both the structural integrity and visual quality of the generated video.
Together, these three mechanisms form a cohesive inference-time guidance framework that achieves robust and precise trajectory control while preserving the generalization ability of VDM priors. Our method is fully training-free and plug-and-play, enabling broad applicability across tasks without model retraining. It is also model-agnostic and readily adapts to backbones such as Wan 2.1 (Wan et al., 2025) and SVD (Blattmann et al., 2023), underscoring its general utility. To validate the effectiveness of the proposed method, we conduct comprehensive experiments on multiple tasks and benchmarks. Results demonstrate consistent improvements in trajectory adherence, geometric consistency, and perceptual quality compared to state-of-the-art (SOTA) baselines. Our main contributions are summarized as follows:
  • We introduce a novel, training-free paradigm for leveraging video generative priors in spatial intelligence tasks, enabling precise and stable 3D/4D trajectory control without retraining or fine-tuning.
  • We design a synergistic inference-time guidance framework integrating Intra-Step Recursive Refinement (IRR) and Flow-Gated Latent Fusion (FLF), achieving accurate trajectory adherence while disentangling motion from content.
  • We propose Dual-Path Self-Corrective Guidance (DSG), a self-referential correction mechanism that enhances spatial alignment and perceptual fidelity without auxiliary networks or retraining.
  • We demonstrate, through extensive experiments on diverse datasets and tasks, that our approach achieves state-of-the-art controllability and visual quality, even compared to training-intensive pipelines.
Figure 2: Overview of our porposed method. Given a single image or video frames, a vision foundation model reconstructs a scene point cloud, which is warped and rendered along a user-specified trajectory to produce a guidance video. The input image (or first frame) is also converted into a textual prompt and latent representation for an image-to-video diffusion model. Trajectory control is injected through a training-free strategy comprising IRR, FLF, and DSG (detailed in Sec. 3.2–3.4), enabling precise control and high-quality synthesis without additional training.
2 Related Works
We review prior work in three areas most relevant to ours: 3D static scene generation, 4D trajectory-controlled video generation, and guidance strategies for generative models.
3D Static Scene Generation. Recent advances in 3D reconstruction (Mildenhall et al., 2020; Kerbl et al., 2023; Song et al., 2024; Gao et al., 2024; Yu et al., 2024d; Müller et al., 2022; Yao et al., 2018) and object-level generation (Poole et al., 2023; Wei et al., 2024; Xiang et al., 2025; Kwak et al., 2024) have achieved strong results, but they lack scene-level priors and struggle with geometric consistency, limiting their scalability. VDM (Blattmann et al., 2023; Wan et al., 2025; Kong et al., 2024) offer richer priors and have thus become a foundation for scene generation. Approaches vary: some decode 3D scenes from a single image via latent traversal (e.g., Wonderland (Liang et al., 2025)), others fine-tune VDMs with depth-based warped inputs, (e.g., ViewCrafter(Yu et al., 2024c) and See3D (Ma et al., 2025a)), and some methods such as MotionCtrl (Wang et al., 2024c) and TrajectoryAttention (Xiao et al., 2025) embed camera parameters to guide view synthesis . More efficient, training-free strategies such as NVS-Solver (You et al., 2025) and ViewExtrapolator (Liu et al., 2024) warp input views and modulate frozen models during sampling. Fine-tuning offers controllability but risks eroding pretrained priors and incurs high cost, while training-free methods retain priors and efficiency but must address geometric coherence. Our work follows the training-free direction, aiming to ensure both view consistency and controllability.
Trajectory-Controlled Dynamic Video Generation. Video synthesis with controllable camera motion generally follows two main paradigms. The first is fine-tuning-based (Ma et al., 2025b; Mou et al., 2024; Yu et al., 2024b; Wang et al., 2024c) , where lightweight adapters $a_{\phi}$ (e.g., LoRA (Hu et al., 2022), ControlNet (Zhang et al., 2023)) are trained on video–trajectory pairs to optimize the standard diffusion (Ho et al., 2020) loss. Examples include ReCamMaster (Bai et al., 2025a), which retrains on 136K annotated videos, TrajectoryCrafter (Yu et al., 2025) with dual-stream conditioning on source videos and 3D point-cloud renders, GCD (Van Hoorick et al., 2024) using synthetic multi-view videos, and DaS (Gu et al., 2025) incorporating 3D tracking signals into a pre-trained diffusion model for multi-type motion control. The second paradigm is warp-and-repaint, which projects source frames with depth to target poses and then repaints the occluded regions (Ma et al., 2025b; Liu et al., 2025; Huang et al., 2025; Tian et al., 2025). While flexible, this approach remains vulnerable to noisy warps that cause flicker or distortion—as in training-free baselines such as NVS-Solver (You et al., 2025), which modulates a frozen video diffusion model using warped views at inference time. In contrast, our method applies a more powerful inference-time guidance mechanism that extracts trajectory cues and directly steers the diffusion process, achieving precise motion-consistency control without any training.
Guidance and Control for Generative Models. A central challenge in diffusion models is how to steer the generative process toward desired outputs. Guidance strategies address this by modifying the sampling trajectory to better satisfy conditioning signals. The most common is Classifier-Free Guidance (CFG) (Ho & Salimans, 2021), which biases generation toward a target condition by blending conditional and unconditional score predictions. While effective, high guidance weights can cause distortions. More advanced methods, such as Auto-Guidance and STG (Karras et al., 2024; Hyung et al., 2025; Xu et al., 2023), use an auxiliary model to anticipate and avoid failure modes, whereas Z-sampling (Bai et al., 2025c) alternates denoising and inversion to refine results mid-generation. For 3D and 4D synthesis, enforcing viewpoint consistency often follows the fine-tuning and warp-and-repaint strategies discussed above. However, the latter remains sensitive to noisy warps (Cai et al., 2024; Wang et al., 2024a), leading to flicker and geometric distortions. To this end, we propose Dual-Path self-corrective guidance, which derives a correction signal from the difference between guided and unguided predictions at each step, enhancing trajectory adherence and stability without retraining or per-scene tuning.
3 Proposed Methods
We address the challenge of balancing controllability, visual fidelity, and generalization when applying video diffusion models (VDMs) to 3D/4D tasks. Our solution is a training-free framework for trajectory-controlled video generation. At its core is an inference-time guidance strategy that steers a pre-trained diffusion model along user-defined trajectories while preserving its intrinsic priors and generative quality. As shown in Fig. 2, the framework integrates three complementary components. Intra-Step Recursive Refinement (IRR) injects trajectory guidance from observed regions at each denoising step, ensuring consistent control throughout sampling (see Sec. 3.2). Flow-Gated Latent Fusion (FLF) refines trajectory injection by decoupling motion and appearance features in the latent space, which prevents content drift and preserves fidelity (see Sec. 3.3). Dual-Path Self-Corrective Guidance (DSG) enhances stability by comparing guided and unguided predictions, using their difference as a corrective signal to suppress artifacts from noisy priors (see Sec. 3.4). Together, these modules unlock the model’s latent 3D/4D awareness and enable fine-grained trajectory control without retraining.
3.1 Preliminaries
Before detailing our method, we introduce the necessary preliminaries: diffusion models, guidance strategies, and trajectory-controlled video synthesis.
3.1.1 Denoising Diffusion Models and Guidance
Diffusion Solvers. Modern generative modeling is dominated by two paradigms: diffusion models (Ho et al., 2020; Song et al., 2020a) and flow-based models (Lipman et al., 2022). Under the SDE view, diffusion models admit a deterministic ODE limit that connects them to flow-based formulations through reparameterization (Gao et al., 2025). The detailed derivation of this equivalence is provided in Appendix A. In this section, we take diffusion models as representative examples. Consider the widely used DDIM sampler (Song et al., 2020a): it recovers the clean sample $\mathbf{x}_{0}$ by reversing the forward noising of a Gaussian prior $\mathbf{x}_{T}$. Given a noise-prediction network $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$, the sampler estimates an intermediate signal $\hat{\mathbf{x}}_{0}$ from the current state $\mathbf{x}_{t}$:
$\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},t)=\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\bar{\alpha}_{t}}},$ (1)
where $\bar{\alpha}_{t}$ denotes cumulative noise attenuation. The term $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},t)$ is a key intermediate variable: at each step, it is the one-step denoised estimate from $\bm{\epsilon}_{\theta}$, evolving from a coarse prediction to a sharp final output. The next sample $\mathbf{x}_{t-1}$ is then obtained by blending $\hat{\mathbf{x}}_{0}$ with the predicted noise $\bm{\epsilon}_{\theta}$:
$\mathbf{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\,\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},t)+\sqrt{1-\bar{\alpha}_{t-1}}\,\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t).$ (2)
Iterating this update from $t=T$ to $t=0$ produces the final sample $\mathbf{x}_{0}$. Our method intervenes at this stage by modifying $\hat{\mathbf{x}}_{0}$ to enforce trajectory control. Notably, other popular solvers, such as UniPC, EDM, and PNDM (Zhao et al., 2023; Karras et al., 2022; Liu et al., 2022a), also compute $\hat{\mathbf{x}}_{0}$ directly or can recover it via a parameterized transformation, so our framework is broadly compatible.
Classifier Free Guidance. To improve fidelity to the condition, CFG (Ho & Salimans, 2021) adjusts the score function during sampling:
$\tilde{\bm{\epsilon}}_{\theta}(\mathbf{x}_{t},t)=\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,\phi)+\omega_{\text{CFG}}\cdot\left[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,\phi)\right],$ (3)
where $\omega_{\text{CFG}}$ is the guidance weight, with $\mathbf{c}$ and $\phi$ denoting the conditional and unconditional branches, respectively. This interpolates conditional and unconditional scores to steer the sampling trajectory. Our approach extends this principle through a self-referential guidance mechanism that dynamically adjusts the guided prediction using the model’s own unguided output at each step.
3.1.2 Trajectory Control via Depth-Based Warping
Our framework uses a warping-and-repainting strategy, embedding geometric cues through depth-based view warping. For reliable depth estimation to support warping into new views, we employ depth prediction networks that takes one or more input images $\{\mathbf{I}_{i}\}_{i=1}^{N}$ and estimates corresponding camera poses and depth maps:
$f:\{\mathbf{I}_{i}\}_{i=1}^{N}\rightarrow\{\mathbf{P}_{i},\mathbf{D}_{i}\},$ (4)
where the depth maps $\mathbf{D}_{i}$ and poses $\mathbf{P}_{i}$ are then used to warp the source views to target poses. Formally, the warping function $\mathcal{W}$ projects source frames $\mathbf{I}_{src}$ with depth $\mathbf{D}_{src}$ from pose $\mathbf{P}_{src}$ to target pose $\mathbf{P}_{tar}$, producing partial target views $\mathbf{I}^{\prime}_{tar}$ and validity masks $\mathbf{M}_{tar}$, which indicate the visible pixels in warped views:
$(\mathbf{I}^{\prime}_{tar},\mathbf{M}_{tar})=\mathcal{W}(\mathbf{I}_{src},\mathbf{D}_{src},\mathbf{P}_{src},\mathbf{P}_{tar}).$ (5)
The warped frames $\mathbf{I}^{\prime}_{tar}$ and masks $\mathbf{M}_{tar}$ provide trajectory-aware observations that serve as the basis for generating videos along arbitrary target poses $\mathbf{P}_{tar}$, though limited to regions visible in source views.
With these preliminaries, we now present our method. Our goal is to use the trajectory control introduced in this section to guide video generation in VDM. To achieve precise and consistent control, we design three modules: Intra-Step Recursive Refinement, Flow-Gated Latent Fusion, and Dual-Path Self-Corrective Guidance. Together, they complement each other to ensure accurate trajectory guidance while preserving high-quality synthesis.
3.2 Intra-Step Recursive Refinement
To enable precise trajectory injection during VDM’s inference processing, we introduce Intra-Step Recursive Refinement (IRR). As noted in Sec. 3.1.1, the denoising process produces an intermediate variable $\hat{\mathbf{x}}_{0}^{(i)}$, a coarse estimate of the final output and the baseline for later steps. Building on this, IRR modifies $\hat{\mathbf{x}}_{0}^{(i)}$ to impose trajectory constraints, ensuring that generation follows the desired path.
IRR operates within the update process of Eq. (1) and Eq. (2). Given the one-step denoised sample $\hat{\mathbf{x}}_{0}^{(i)}$ from Eq. (1), we fuse it with the trajectory latent $\mathbf{Z}_{\text{traj}}$, obtained by encoding the warped frames from Eq. (5) into the latent space. We then add small Gaussian noise $\bm{\epsilon}$ to obtain a modified latent $\mathbf{x}_{t_{i}}^{\prime}$:
$\mathbf{x}_{t_{i}}^{\prime}=(1-w(\sigma))\,\mathbf{F}(\hat{\mathbf{x}}_{0}^{(i)},\mathbf{Z}_{\text{traj}})+w(\sigma)\,\cdot\bm{\epsilon},$ (6)
where $\bm{\epsilon}=\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is the initial Gaussian noise used in sampling. To return the fused result $\mathbf{x}_{t_{i}}^{\prime}$ into the denoising process while injecting trajectory information, we reintroduce noise $\bm{\epsilon}$. Here $0<w(\cdot)<1$ is a user-defined weighting factor derived from the noise schedule $\sigma$, which controls the strength of the reintroduced noise. $\mathbf{F}(\cdot)$ is a mask-based fusion function defined as $\mathbf{F}(\hat{\mathbf{x}}_{0}^{(i)},\mathbf{Z}_{\text{traj}})=\mathbf{M}\cdot\mathbf{Z}_{\text{traj}}+(1-\mathbf{M})\cdot\hat{\mathbf{x}}_{0}^{(i)},$ where $\mathbf{M}$ is the binary validity mask from Eq. (5). It copies observable warped content from $\mathbf{Z}_{\text{traj}}$ into the corresponding locations of $\hat{\mathbf{x}}_{0}^{(i)}$, while leaving unobserved regions unchanged.
In summary, IRR embeds a micro predictor–corrector at each denoising step. By updating $\hat{\mathbf{x}}_{0}^{(i)}$ with explicit trajectory cues, it continually corrects the sampling path and ensures that synthesis follows the target trajectory precisely.
3.3 Flow-Gated Latent Fusion
In the IRR process, directly overwriting all latent channels with trajectory information often degrades visual quality. This is because VAE latents tend to encode different types of information: some channels mainly capture appearance, while others relate more to motion. Using Eq. (6) to replace all channels indiscriminately introduces noise into appearance-dominant ones. To avoid this, we introduce Flow-Gated Latent Fusion (FLF), which decouples motion and appearance features in the latent space and selectively injects trajectory guidance into motion-relevant channels.
To separate motion from appearance and identify motion-relevant channels, we introduce a flow-based scoring scheme. Optical flow captures pixel-wise motion between frames and is widely used to describe temporal dynamics. By comparing predicted and reference flows, we estimate how strongly each latent channel encodes motion.
Given the predicted latent $\hat{\mathbf{x}}_{0}^{(i)}$ from IRR and the reference trajectory latent $\mathbf{Z}_{\text{traj}}$, we compute per-channel flows using the Farnebäck algorithm. For each channel $c$, we obtain predicted flow $\mathcal{F}^{(c)}_{\text{pred}}(x,y,\tau)=(u^{(c)}_{\text{pred}}(x,y,\tau),v^{(c)}_{\text{pred}}(x,y,\tau))$ and reference ground truth (GT) flow $\mathcal{F}^{(c)}_{\text{gt}}(x,y,\tau)=(u^{(c)}_{\text{gt}}(x,y,\tau),v^{(c)}_{\text{gt}}(x,y,\tau))$, restricted to valid regions $\mathbf{M}^{(c)}(x,y,\tau)$ from Eq. (5). Here $(x,y)$ are pixel coordinates, $\tau$ is the time index, and $u_{*},v_{*}$ are the horizontal and vertical components.
To evaluate motion relevance, we design a normalized score that combines three masked metrics: Masked End-point Error (M-EPE), Masked Angular Error (M-AE), and Outlier Percentage (Fl-all). M-EPE is the average distance between the vectors of the predicted flow and GT flow; smaller values mean better alignment. M-AE measures the angular difference between flows; smaller values mean more accurate directions. Fl-all is the percentage of pixels with large deviations; lower values mean fewer outliers and higher reliability. Together, these metrics provide a robust estimate of each channel’s motion relevance. The formulas are given below:
M-EPE: Given the set $\Omega$ of all valid pixels $(x,y,\tau)$ where $\mathbf{M}^{(c)}=1$, we compute the per-channel $L_{2}$ distance between predicted and GT flows and normalize by the number of valid pixels $|\Omega|$:
$\text{M-EPE}_{c}=\frac{1}{|\Omega|}\sum_{(x,y,\tau)\in\Omega}\left\|\mathcal{F}^{(c)}_{\text{pred}}(x,y,\tau)-\mathcal{F}^{(c)}_{\text{gt}}(x,y,\tau)\right\|_{2}.$ (7)
M-AE: Similar to M-EPE, for each valid pixel in $\Omega$ we compute the cosine similarity between $\mathcal{F}^{(c)}_{\text{pred}}$ and $\mathcal{F}^{(c)}_{\text{gt}}$, and use it to measure angular error:
$\text{M-AE}_{c}=\frac{1}{|\Omega|}\sum_{(x,y,\tau)\in\Omega}\arccos\left(\frac{\mathcal{F}^{(c)}_{pred}(x,y,\tau)\cdot\mathcal{F}^{(c)}_{gt}(x,y,\tau)}{\|\mathcal{F}^{(c)}_{pred}(x,y,\tau)\|\cdot\|\mathcal{F}^{(c)}_{gt}(x,y,\tau)\|}\right).$ (8)
Fl-all: A valid pixel in $\Omega$ is an outlier if its M-EPE exceeds $3$ or its relative error exceeds $5\%$. Fl-all (denoted as $\text{F}_{c}$) is the percentage of outliers among valid pixels; lower values mean higher reliability.
To unify the scales of the three metrics and simplify score computation, we normalize $\text{M-EPE}_{c},\text{M-AE}_{c},\text{F}_{c}$ to $[0,1]$, denoted as $\text{E}^{\text{Norm}}_{c},\text{A}^{\text{Norm}}_{c},$ and $\text{F}^{\text{Norm}}_{c}$, respectively:
$\text{E}^{\text{Norm}}_{c}=\min(\text{M-EPE}_{c}/n_{1},1),\quad\text{A}^{\text{Norm}}_{c}=\min(\text{M-AE}_{c}/n_{2},1),\quad\text{F}^{\text{Norm}}_{c}=\min(\text{F}_{c}/n_{3},1),$ (9)
where $n_{1},n_{2},n_{3}$ are normalization factors. We then define the final flow score for each channel as $S$, which directly reflects the consistency between the predicted and GT flows:
$S^{(c)}=\gamma_{1}\!\left(1-\text{E}^{\text{Norm}}_{c}\right)+\gamma_{2}\!\left(1-\text{A}^{\text{Norm}}_{c}\right)+\gamma_{3}\!\left(1-\text{F}^{\text{Norm}}_{c}\right),$ (10)
where $\gamma_{1}+\gamma_{2}+\gamma_{3}=1$. The flow score $S^{(c)}$ measures the motion relevance of channel $c$. A higher score means its predicted flow is closer to the ground truth and thus encodes more motion. A lower score indicates misalignment and suggests that the channel encodes more non-motion information, such as appearance. This score enables us to decouple motion and appearance channels in the latent space. To select motion-relevant channels, we set a dynamic threshold:
$\delta^{(i)}=\mu_{S}^{(i)}-\lambda^{(i)}\sigma_{S}^{(i)},$ (11)
where $\mu_{S}^{(i)}$ and $\sigma_{S}^{(i)}$ are the mean and standard deviation of $\{S^{(i,c)}\}$ at step $i$. Channels with scores above $\delta^{(i)}$ are regarded as motion-relevant and only these channels receive trajectory guidance. The selection sets are:
$\mathcal{C}_{\text{sel}}^{(i)}=\{\,c\mid S^{(i,c)}\geq\delta^{(i)}\,\}.$
Only these channels are updated with trajectory latents $\mathbf{Z}_{\text{traj}}$. The parameter $0<\lambda^{(i)}<1$ controls the strictness of selection and decreases during denoising. Early steps therefore replace more channels for stronger trajectory adherence, while later steps replace fewer channels to preserve appearance details. Finally, the latent update rule for $\hat{\mathbf{x}}_{0}^{(i)}$ is:
$\mathbf{FLF}(\hat{\mathbf{x}}_{0}^{(i)},\mathbf{Z}_{\text{traj}})^{(c)}=\begin{cases}\mathbf{M}^{(c)}\cdot\mathbf{Z}_{\text{traj}}^{(c)}+\left(1-\mathbf{M}^{(c)}\right)\cdot\hat{\mathbf{x}}_{0}^{(i,c)},&c\in\mathcal{C}_{\text{sel}}^{(i)},\\[2.0pt] \hat{\mathbf{x}}_{0}^{(i,c)},&\text{otherwise}.\end{cases}$ (12)
where $\mathbf{FLF}(\cdot)$ denotes the flow-score–based selective fusion. We use it to replace $\mathbf{F}(\cdot)$ in Eq. (6), which overwrote all channels without distinction. The updated fusion rule is:
$\mathbf{x}_{t_{i}}^{\prime}=(1-w(\sigma))\,\mathbf{FLF}(\hat{\mathbf{x}}_{0}^{(i)},\mathbf{Z}_{\text{traj}})+w(\sigma)\,\cdot\bm{\epsilon},$ (13)
In summary, FLF enables fine-grained trajectory injection while preserving model priors and synthesis quality. Unlike Restart Sampling (Xu et al., 2023), which restarts the schedule, or ViewExtrapolator (Liu et al., 2024), which injects trajectory data directly into $\hat{\mathbf{x}}_{0}^{(i)}$, our method uses IRR to apply fine-grained guidance at every timestep. Combined with FLF, it separates motion from appearance and delivers precise, high-quality trajectory control at inference.
3.4 Dual-Path Self-Corrective Guidance
Trajectory latents $\mathbf{Z}_{\text{traj}}$ obtained by warping along the target motion often suffer from distortions due to depth errors, occlusions, or misalignments. Such artifacts are common in warp-based methods and degrade synthesis quality. To address this, we draw inspiration from CFG (Ho & Salimans, 2021), which interpolates between conditional and unconditional predictions. Instead of using conditioning, we exploit the discrepancy between two denoising paths in IRR to obtain a more reliable direction. Based on this idea, we propose Dual-Path Self-Corrective Guidance (DSG). At each iteration, IRR produces two velocity fields: the unguided $\mathbf{v}_{t_{i}}^{\text{ori}}$ from the original latent $\mathbf{x}_{t_{i}}$, and the guided $\mathbf{v}_{t_{i}}^{\text{traj}}$ from the corrected latent $\mathbf{x}^{\prime}_{t_{i}}$ obtained by injecting trajectory cues. DSG leverages their difference to form a correction signal that steers denoising toward a higher-quality path.
At each step $t_{i}$, IRR and FLF generate two velocity fields: $\mathbf{v}_{t_{i}}^{\text{ori}}$ and $\mathbf{v}_{t_{i}}^{\text{traj}}$. The guided velocity $\mathbf{v}_{t_{i}}^{\text{traj}}$ enforces trajectory adherence but may add noise from imperfect warping, while the unguided velocity $\mathbf{v}_{t_{i}}^{\text{ori}}$ stays on the data manifold with higher fidelity but ignores trajectory cues. To exploit these two directions for a better denoising path, we propose DSG, which combines them to balance control and fidelity. Unlike CFG (Eq. (3)) (Ho & Salimans, 2021), which interpolates unconditional and conditional predictions, DSG uses two trajectory-specific complementary branches designed for motion control. The corrected velocity $\mathbf{v}_{t_{i}}^{\text{corr}}$ is then computed as:
$\mathbf{v}_{t_{i}}^{\text{corr}}=\mathbf{v}_{t_{i}}^{\text{traj}}+\rho\cdot\beta_{t_{i}}\bigl(\mathbf{v}_{t_{i}}^{\text{traj}}-\mu\cdot\alpha_{t_{i}}\cdot\mathbf{v}_{t_{i}}^{\text{ori}}\bigr),$ (14)
where $\rho$ controls guidance strength, $\mu$ normalizes magnitudes, $\alpha_{t_{i}}=(\mathbf{v}_{t_{i}}^{\text{traj}}\cdot\mathbf{v}_{t_{i}}^{\text{ori}})/(\lVert\mathbf{v}_{t_{i}}^{\text{traj}}\rVert\cdot\lVert\mathbf{v}_{t_{i}}^{\text{ori}}\rVert)$ is the cosine similarity between guided and unguided velocities, and $\beta_{t_{i}}=\sin(\arccos(\alpha_{t_{i}}))$ is the corresponding sine similarity. This adaptive weighting amplifies corrections when the two paths diverge (lower $\alpha_{t_{i}}$, higher $\beta_{t_{i}}$), pulling the result toward the guided direction, and reduces corrections when they agree (higher $\alpha_{t_{i}}$, lower $\beta_{t_{i}}$), preserving the model’s natural prediction.
In score-based (Lipman et al., 2022) terms, $\mathbf{v}_{t_{i}}^{\text{ori}}$ follows the pretrained data manifold, while $\mathbf{v}_{t_{i}}^{\text{traj}}$ may deviate due to injected trajectory signals. DSG applies cosine-weighted interpolation between the two, dynamically combining them to retain trajectory guidance while projecting the velocity back toward the manifold. This improves generation quality by suppressing off-manifold drift. As a result, $\mathbf{v}_{t_{i}}^{\text{corr}}$ drives the sample along the desired motion path while preserving visual fidelity to the model priors. Subsequent experiments confirm the effectiveness of this mechanism.
4 Experiments
In this section, we present a comprehensive evaluation of our proposed training-free framework. We first outline the implementation details in Sec. 4.1. Subsequently, we demonstrate the performance of our method on 3D scene generation and 4D trajectory control in Sec. 4.2 and Sec. 4.3, respectively. Finally, we conduct a series of ablation studies in Sec. 4.4 to validate the effectiveness of each component of our approach.
4.1 Implementation Details
Our framework is a training-free, inference-time optimization method that steers pre-trained video diffusion models for precise camera control without additional training or fine-tuning. It introduces no significant computational overhead beyond the base model’s inference requirements.
[…]
Our pipeline follows a warp-and-repaint design. For warping, we test several depth estimation models, such as VGGT (Wang et al., 2025b), UniDepth (Piccinelli et al., 2024), Mega-SaM (Li et al., 2025), and DepthCrafter (Hu et al., 2025). The method adapts well to all, benefiting from the strong world priors of the underlying video model.
Test Datasets and Metrics. For single-view 3D scene generation, we adopt the widely used LLFF (Mildenhall et al., 2019), Tanks and Temples (Knapitsch et al., 2017), and MipNeRF 360 (Barron et al., 2022) datasets, selecting half of the scenes from each for evaluation. We also test on diverse internet, real-world, and AI-generated images. Perceptual quality is assessed using FID (Heusel et al., 2017) and $\mathrm{CLIP}_{\mathrm{sim}}$ (Radford et al., 2021) similarity. For 4D trajectory control, we compare with SOTA methods on their respective benchmarks, which include challenging real-world video clips with varied camera trajectories.
Performance is measured with FVD (Unterthiner et al., 2018) for temporal quality and $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$ for source–target consistency. In addition, we assess camera trajectory accuracy using three standard metrics: Absolute Trajectory Error (ATE), which captures the global alignment between estimated and reference paths; Relative Pose Error – Translation (RPE-T), which reflects short-term consistency of translational motion; and Relative Pose Error – Rotation (RPE-R), which quantifies local orientation accuracy. Together, these metrics provide a comprehensive evaluation of both global fidelity and local smoothness of camera trajectories.
4.2 3D Scene Generation
Figure 3: Qualitative comparison of novel view synthesis from a single input image. The first row shows the input frame and its depth-based warped views, where disoccluded regions appear as holes. Compared to existing SOTA methods, our approach produces more consistent scene content under novel viewpoints, with improved image detail, trajectory accuracy, and structural plausibility.
We compare our method with SOTA approaches for novel view synthesis, which can be grouped into two categories. The first includes methods requiring model-specific training or fine-tuning, such as ViewCrafter (Yu et al., 2024c), which trains a VDM on iteratively refined point clouds; TrajectoryCrafter (Yu et al., 2025), which learns motion control via dedicated training; TrajectoryAttention (Xiao et al., 2025), which injects motion through a specialized attention module; and See3D(Ma et al., 2025a), a video-based model for scene reconstruction. The second category comprises training-free methods, such as NVS-Solver (You et al., 2025), which modulates a pre-trained model using warped views, and ViewExtrapolator (Liu et al., 2024), which employs an annealed guidance strategy to improve trajectory adherence.
Table 1: Quantitative results on static 3D and dynamic 4D benchmarks. We evaluate on public datasets such as LLFF and MipNeRF360, as well as Internet images and videos. For static scenes we report $\mathrm{CLIP}_{\mathrm{sim}}$ and FID, and for dynamic scenes $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$ and FVD. Metric details are provided in Appendix B. Our method consistently surpasses SOTA baselines in generation quality.
Static Dynamic
FID $\downarrow$ $\mathrm{CLIP}_{\mathrm{sim}}\uparrow$ FVD $\downarrow$ $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}\uparrow$
See3D (Ma et al., 2025a)
123.26 0.941 - -
ViewCrafter (Yu et al., 2024c)
117.50 0.930 - -
ViewExtrapolator (Liu et al., 2024)
125.50 0.930 108.48 0.913
TrajectoryAttention (Xiao et al., 2025)
122.37 0.920 106.94 0.911
TrajectoryCrafter (Yu et al., 2025)
111.49 0.910 97.31 0.923
NVS-Solver (You et al., 2025)
118.64 0.937 - -
WorldForge (Ours)
96.08 0.948 93.17 0.938
Table 2: Quantitative comparison of camera-trajectory accuracy on static and dynamic scenes. We evaluate against SOTA methods using ATE, RPE-T, and RPE-R, lower is better. Across both settings, our method achieves the best or second-best results on all metrics. Metric definitions and computation details are provided in Appendix B.
Static Dynamic
ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$ ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$
See3D (Ma et al., 2025a)
0.091 0.089 0.250 - - -
ViewCrafter (Yu et al., 2024c)
0.236 0.315 0.728 - - -
ViewExtrapolator (Liu et al., 2024)
0.183 0.260 0.882 1.040 1.208 4.750
TrajectoryAttention (Xiao et al., 2025)
0.159 0.238 0.532 0.605 1.238 3.560
TrajectoryCrafter (Yu et al., 2025)
0.090 0.152 0.267 0.431 1.078 8.950
NVS-Solver (You et al., 2025)
0.224 0.268 1.056 - - -
WorldForge (Ours)
0.077 0.086 0.221 0.527 0.826 2.690
[…]
Frames Resolution
Inference Time
Base Video
Model
Training-
Free
See3D (Ma et al., 2025a)
25 $576\times 1024$ 1.7 Custom
ViewCrafter (Yu et al., 2024c)
25 $576\times 1024$ 1.8 Custom
ViewExtrapolator (Liu et al., 2024)
25 $576\times 1024$ 1.6 SVD
TrajectoryAttention (Xiao et al., 2025)
25 $76\times 1024 5.5$ SVD
TrajectoryCrafter (Yu et al., 2025)
25 $384\times 672$ 1.7 CogVideoX
NVS-Solver (You et al., 2025)
25 $576\times 1024$ 9.3 SVD
ReCamMaster (Bai et al., 2025a)
81 $480\times 832$ 14.6 Wan 2.1 T2V
WorldForge (Ours, 720P)
25 $1280\times 720$ 17.3 Wan 2.1 I2V
WorldForge (Ours, 480P)
25 $480\times 832$ 6.8 Wan 2.1 I2V
WorldForge (Ours, on SVD)
25 $576\times 1024$ 1.3 SVD
We evaluate all methods under consistent settings with the same input image, camera trajectory, depth, and target scenes across LLFF (Mildenhall et al., 2019), MipNeRF-360 (Barron et al., 2022), and Tanks-and-Temples (Knapitsch et al., 2017). Results (Figs.3, Table1) show that our inference-time guidance consistently outperforms both training-based and training-free baselines. Moreover, we further assess human-centric scenes, which are more sensitive to visual consistency and content quality. As shown in Fig. 4, prior methods often fail in such cases. In contrast, our approach, supported by the proposed guidance strategy, preserves model priors and faithfully follows the target trajectory. This offers an effective paradigm for reconciling controllability with generalization and high fidelity. To independently validate the effect of our guidance strategy, we further evaluate on the lighter U-Net-based SVD model (Blattmann et al., 2023); even in this setting, our method attains superior visual quality.
Figure 4: Static 3D generation on human-centric scenes. Existing methods struggle, particularly with motion-prone shots (left) and portrait close-ups (right). On the left, baselines introduce artifacts and unintended motion. On the right, most fail to produce plausible results; TrajectoryCrafter (Yu et al., 2025) recovers coarse structure but lacks detail and visual appeal. In contrast, our method maintains scene stationarity under trajectory guidance and produces natural, faithful renderings, achieving both precise control and high perceptual quality.
Notably, our framework can synthesize high-quality, photorealistic, and structurally consistent $360^{\circ}$ views from a single input without relying on panoramic intermediates (Fig. 5). This capability extends to complex outdoor and open-world scenes, pushing the limits of single-view novel view synthesis.
Figure 5: 360° orbit views from a single real-world outdoor image. With precise trajectory control and realistic rendering, our method overcomes the viewpoint limitation of single-image generation and produces ultra-wide views of complex real scenes. Unlike panorama-based approaches, it directly supports object-centric trajectories and achieves higher visual quality.
4.3 4D Trajectory Control
For dynamic trajectory control, we compare against leading video-to-video models. These include ReCamMaster (Bai et al., 2025a), which employs a sophisticated conditioning mechanism trained on large-scale synthetic data, TrajectoryCrafter (Yu et al., 2025), with its dual-stream diffusion architecture, TrajectoryAttention (Xiao et al., 2025), which processes motion through an additional attention module, and ViewExtrapolator (Liu et al., 2024).
Following the protocols of prior works, we evaluate on diverse and challenging video clips, including arcs, dolly zooms, and composite paths. Our training-free method consistently delivers higher visual fidelity, better trajectory alignment, and more coherent scene completion. As shown in Table 1 and Fig. 6, it matches or surpasses state-of-the-art methods that require costly training, demonstrating the strength of leveraging latent 3D priors in existing models. Notably, our approach excels at plausibly reconstructing unseen regions, while other methods often produce distortions or implausible content.
Figure 6: Comparison of 4D trajectory-controlled re-rendering. Baselines often produce implausible artifacts (e.g., flattened faces, floating heads), reflecting limited use of pretrained priors. Our inference-time guidance leverages these latent world priors to re-render realistic, high-quality content along the target trajectory. We compare against state-of-the-art baselines under identical inputs; for ReCamMaster (text-controlled), parameters are adjusted to match the target path.
Beyond benchmarks, the framework supports various post-production tasks. It can stabilize videos by smoothing camera motion and control camera paths to enable localized super-resolution or outpainting. In addition, by specifying masked regions, it supports creative video content edits such as object addition or removal, subject replacement, and try-on effects (see Fig. 7). These capabilities highlight its versatility for real-world video re-rendering.
Figure 7: Other video effects enabled by our method. Beyond video re-cam, our flexible depth-based warping also supports various video editing operations, such as freezing the camera, stabilizing video, and editing video content. These extensions further broaden the practical scope of our approach.
4.4 Ablation Experiments
We conducted ablation studies to assess the contribution of each component in our framework.
Figure 8: Ablation of the proposed components. IRR enables trajectory injection; without it, the model defaults to prompt-only free generation, and FLF/DSG cannot be applied. FLF decouples trajectory cues from noisy content; removing it introduces noise from warped frames. DSG guides sampling toward high-quality, trajectory-consistent results; without it, detail and plausibility drop. The full model achieves the best fidelity and control, demonstrating their complementary effects.
Component Analysis. We examined the effects of removing Intra-Step Recursive Refinement (IRR), Flow-Gated Latent Fusion (FLF), and Dual-Path Self-Corrective Guidance (DSG). As shown in Fig. 8 without IRR, trajectory guidance cannot be injected during inference, and the model fails to follow the target path. Removing FLF, i.e., not separating motion and appearance channels, damages the model’s priors and produces unnatural results. Without DSG, noise from warped trajectories propagates into the generation, causing low quality and artifacts. The full model achieves the best results, confirming that these components work synergistically to ensure robust and precise control.
Base Model. We replaced the Wan2.1 (Wan et al., 2025) model with the U-Net-based SVD (Blattmann et al., 2023) model to verify model-agnosticism and transferability (See in Fig. 9). Fair comparisons with other SVD-based methods show that our approach integrates seamlessly into pre-trained SVD. Although limited by SVD’s weaker priors, our method still reaches state-of-the-art performance within the same model. This demonstrates that our guidance is architecture-independent and suggests even greater potential with stronger future base models.
Depth Estimation Model. As shown in Fig. 10 We tested VGGT (Wang et al., 2025b), Mega-SaM (Li et al., 2025), UniDepth (Piccinelli et al., 2024) and DepthCrafter (Hu et al., 2025) for depth-based warping. Each shows distinct strengths and weaknesses, yet our method remains effective across all. Thanks to the generative model’s strong 3D prior, many warping artifacts (e.g., tearing, stretching) are automatically corrected during refinement. This robustness enables plug-and-play integration with various SOTA depth estimation techniques without sacrificing performance.
Figure 9: Ablation across different VDMs. To rule out the influence of the intrinsic performance advantage of the VDM (Wan2.1 (Wan et al., 2025)) and to verify the method’s transferability, we port the proposed guidance to a compact U-Net–based SVD model (Blattmann et al., 2023) and compare against SVD-based SOTA baselines. Experiments show that the guidance transfers seamlessly, makes the native SVD controllable, and achieves SOTA performance in content quality, structural plausibility, and trajectory consistency.
Figure 10: Depth estimation models ablation. Our method leverages the inherent world knowledge of VDMs to correct errors and fill missing regions even under challenging inputs (left). This strong self-correction ability ensures broad compatibility with different depth estimators (right). Despite variations or noise in depth-based warping, it reliably compensates through learned priors and produces realistic, high-quality results.
5 Conclusion and Limitation
We present WorldForge, a training-free framework for trajectory-controllable generation in both static 3D scenes and dynamic 4D scenes. Our method tackles the persistent challenge in controllable video generation of balancing high visual quality, strong generalization, and precise controllability. At its core is a unified guidance strategy—Intra-Step Recursive Refinement (IRR), Flow-Gated Latent Fusion (FLF), and Dual-Path Self-Corrective Guidance (DSG)—that operates entirely at inference time. By decoupling motion from appearance features and correcting trajectory drift caused by structural noise, it injects fine-grained predefined trajectory constraints while preserving the valuable prior world knowledge embedded in the base model. Extensive experiments demonstrate state-of-the-art performance on both 3D and 4D trajectory generation tasks, offering a new paradigm for exploring spatial intelligence and emergent world models in large-scale generative systems.
While our framework can repair many distortions introduced by depth-based warping, it may fail under extremely poor depth estimations (e.g., completely flattened subjects or severe foreground–background entanglement). Moreover, due to the global nature of our guidance, control over small objects or fine details remains limited. Future work will explore integrating fine-grained control mechanisms and applying our approach to more powerful generative models.
Appendix A Proof of the equivalence between diffusion and flow models
We consider Flow Matching (Lipman et al., 2022; Liu et al., 2022b) as a special case of diffusion modeling (Kingma & Gao, 2023; Gao et al., 2025). In the following, we will first outline the formulation of diffusion models and then substitute the specific parameterization of Flow Matching to demonstrate their compatibility.
Given a random variable $\mathbf{x}_{0}$ drawn from an unknown data distribution $q_{0}(\mathbf{x}_{0})$, a Diffusion Probabilistic Model (DPM) (Ho et al., 2020; Song et al., 2020b; Lu et al., 2022) defines a forward process that gradually transforms the data into a simple prior distribution, typically a Gaussian distribution. The conditional distribution of the noised variable $\mathbf{x}_{t}$ at time $t$ given the initial data $\mathbf{x}_{0}$ is defined as a Gaussian transition kernel (Kingma et al., 2021):
$q_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t}|\alpha_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}).$ (15)
Equivalently, a sample $\mathbf{x}_{t}$ at any time $t\in[0,T]$ can be expressed through a reparameterization (Kingma et al., 2021; Gao et al., 2025):
$\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\bm{\epsilon},~\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (16)
Here, $\alpha_{t}$ and $\sigma_{t}$ are scalar functions of time, known as the noise schedule, that control the signal-to-noise ratio. Typically, $\alpha_{t}$ decreases over time while $\sigma_{t}$ increases, satisfying a condition such as $\alpha_{t}^{2}+\sigma_{t}^{2}=1$ in Variance Preserving (VP) SDEs (Ho et al., 2020; Song et al., 2020b). Kingma et al. (2021) proves that the following stochastic differential equation (SDE) has the same transition distribution in Eq. (15) for any $t\in[0,T]$:
$\mathrm{d}\mathbf{x}_{t}=f(t)\mathbf{x}_{t}\mathrm{d}t+g(t)\mathrm{d}\mathbf{w}_{t},~\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0}),$ (17)
where $\mathbf{w}_{t}$ is a standard Wiener process. The drift coefficient $f(t)$ and the diffusion coefficient $g(t)$ can be derived using schedule parameters $\alpha_{t}$ and $\sigma_{t}$ (Kingma et al., 2021):
$f(t)=\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t},\quad g^{2}(t)=\frac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{d}t}-2\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}\sigma_{t}^{2}.$ (18)
The generative process of diffusion models involves reversing this forward process. This can be achieved via a corresponding reverse-time SDE (Song et al., 2020b). For more efficient generation, one can utilize the associated probability flow ordinary differential equation (PF-ODE), which shares the same marginal distributions as at each time $t$ as that of the SDE (Song et al., 2020b). This PF-ODE is given by:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(t)\mathbf{x}_{t}-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}).$ (19)
By relating the score function $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ to the noise term via $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\approx-\frac{\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)}{\sigma_{t}}$, where $\bm{\epsilon}_{\theta}$ is a neural network trained to predict the noise, the ODE becomes (Karras et al., 2022; Zhao et al., 2023):
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(t)\mathbf{x}_{t}+\frac{g^{2}(t)}{2\sigma_{t}}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t).$ (20)
Now, let us consider the forward process in Flow Matching (Lipman et al., 2022; Liu et al., 2022b). The path from a data point $\mathbf{x}_{0}$ to a noise sample $\bm{\epsilon}$ is defined by a simple linear interpolation:
$\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+t\cdot\bm{\epsilon},~\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (21)
where $t\in[0,1]$. By comparing Eq. (21) with the general form of the diffusion forward process in Eq. (16), we can establish a direct correspondence by setting the diffusion schedule parameters as:
$\alpha_{t}=1-t\quad\text{and}\quad\sigma_{t}=t.$
Substituting this specific parameterization into the definitions for $f(t)$ and $g(t)$ in Eq. (18), we derive the corresponding coefficients for this Flow Matching SDE:
$f_{\mathrm{FM}}(t)=\frac{\mathrm{d}\log(1-t)}{\mathrm{d}t}=\frac{-1}{1-t},$ (22)
$g^{2}_{\mathrm{FM}}(t)=\frac{\mathrm{d}(t^{2})}{\mathrm{d}t}-2\frac{-1}{1-t}t^{2}=\frac{2t}{1-t}.$ (23)
Next, we insert these specific coefficients $f_{\mathrm{FM}}(t)$ and $g^{2}_{\mathrm{FM}}(t)$ into the PF-ODE formulation from Eq. (20). To analyze the underlying dynamics, we consider the ideal case where the score is perfectly known, which is equivalent to replacing the model prediction $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ with the ground-truth noise $\bm{\epsilon}$. This yields:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$ $=f_{\mathrm{FM}}(t)\mathbf{x}_{t}+\frac{g_{\mathrm{FM}}^{2}(t)}{2\sigma_{t}}\bm{\epsilon}$
$=\frac{-1}{1-t}\mathbf{x}_{t}+\frac{2t}{2t\cdot(1-t)}\bm{\epsilon}$
$=\frac{\bm{\epsilon}-\mathbf{x}_{t}}{1-t}$
$=\frac{\bm{\epsilon}-[(1-t)\mathbf{x}_{0}+t\cdot\bm{\epsilon}]}{1-t}$
$=\frac{(1-t)\bm{\epsilon}-(1-t)\mathbf{x}_{0}}{1-t}$
$=\bm{\epsilon}-\mathbf{x}_{0}.$ (24)
This resultant vector field, $\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=\bm{\epsilon}-\mathbf{x}_{0}$, is precisely the time derivative of the Flow Matching path defined in Eq. (21). This equivalence demonstrates that the process prescribed by Flow Matching is a specific instance of the diffusion models, corresponding to the linear noise schedule $\alpha_{t}=1-t$ and $\sigma_{t}=t$. Therefore, Flow Matching can be formally viewed as a subset of the broader diffusion modeling framework (Kingma & Gao, 2023; Gao et al., 2025).
Appendix B Evaluation Metrics
We employ seven complementary metrics to comprehensively evaluate video generation quality: FID and $\mathrm{CLIP}_{\mathrm{sim}}$ similarity for static scenes, FVD and $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$ for dynamic scenes, and ATE, RPE-T, and RPE-R for camera trajectory consistency. These metrics provide objective quantitative assessment across multiple dimensions including image realism, semantic consistency, temporal coherence, and camera motion fidelity.
B.1 Static Scene Evaluation
Fréchet Inception Distance (FID). FID (Heusel et al., 2017) measures image generation quality by comparing the distribution of real and generated images in the Inception-V3 feature space. We use an ImageNet-pretrained Inception-V3 model and extract 2048-dimensional features from the pool3 layer. The FID score is computed as:
$\text{FID}=\|\mu_{r}-\mu_{g}\|^{2}+\text{Tr}(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{1/2})$ (25)
where $\mu_{r}$ and $\mu_{g}$ are the mean vectors of real and generated image features, and $\Sigma_{r}$ and $\Sigma_{g}$ are the corresponding covariance matrices.
CLIP Similarity. CLIP similarity (Radford et al., 2021) evaluates the semantic similarity between generated and real images using vision-language pre-trained representations. We employ the CLIP ViT-B/32 model trained on 400 million image-text pairs. The similarity score is calculated as:
$\text{CLIP}_{\text{sim}}=\frac{1}{N}\sum_{i=1}^{N}\cos(f_{r,i},f_{g,i})$ (26)
where $f_{r,i}$ and $f_{g,i}$ are the L2-normalized 512-dimensional CLIP features of the $i$-th real and generated image pair.
B.2 Dynamic Scene Evaluation
Fréchet Video Distance (FVD). FVD (Unterthiner et al., 2018) extends FID to the video domain by measuring distribution differences in spatio-temporal feature space. For computational efficiency, we implement a simplified version that treats video frame sequences as image collections and applies Inception-V3 frame-wise feature extraction. The FVD score follows the same Fréchet distance formula as FID but operates on video frame features.
Video CLIP Similarity ($\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$). $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$ extends CLIP similarity to the temporal domain by computing frame-level semantic consistency between generated and real videos. The score is calculated as:
$\mathrm{CLIP\text{-}V}_{\mathrm{sim}}=\frac{1}{M}\sum_{j=1}^{M}\left[\frac{1}{T_{j}}\sum_{t=1}^{T_{j}}\cos(f_{r,j,t},f_{g,j,t})\right]$ (27)
where $M$ is the number of video pairs, $T_{j}$ is the frame count of the $j$-th video pair, and $f_{r,j,t}$, $f_{g,j,t}$ are the CLIP features of the $t$-th frame in the $j$-th video pair.
B.3 Camera Trajectory Evaluation
Absolute Trajectory Error (ATE). ATE measures the global consistency between estimated and reference camera trajectories by computing Euclidean distances between corresponding camera positions. The ATE for each timestamp is computed as:
$\text{ATE}_{i}=\|\mathbf{t}_{\text{ref},i}-\mathbf{t}_{\text{est},i}\|_{2}$ (28)
where $\mathbf{t}_{\text{ref},i}$ and $\mathbf{t}_{\text{est},i}$ are the 3D position vectors of the reference and estimated trajectories at timestamp $i$, respectively. The root mean square error over all $n$ trajectory points is:
$\text{ATE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\text{ATE}_{i}^{2}}$ (29)
Relative Pose Error - Translation (RPE-T). RPE-T evaluates the local accuracy of camera translation between consecutive frames, reflecting short-term trajectory consistency. For consecutive frames, the relative translation error is:
$\text{RPE-T}_{i}=\|\Delta\mathbf{t}_{\text{ref},i}-\Delta\mathbf{t}_{\text{est},i}\|_{2}$ (30)
where $\Delta\mathbf{t}_{\text{ref},i}=\mathbf{t}_{\text{ref},i+1}-\mathbf{t}_{\text{ref},i}$ and $\Delta\mathbf{t}_{\text{est},i}=\mathbf{t}_{\text{est},i+1}-\mathbf{t}_{\text{est},i}$. The RMSE over all n1n-1 consecutive frame pairs is:
$\text{RPE-T}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n-1}\text{RPE-T}_{i}^{2}}$ (31)
Relative Pose Error - Rotation (RPE-R). RPE-R assesses the accuracy of camera orientation changes between consecutive frames. For rotation matrices $\mathbf{R}_{\text{ref},i}$ and $\mathbf{R}_{\text{est},i}$, the angular difference in degrees is computed as:
$\text{RPE-R}_{i}=\arccos\left(\frac{\text{trace}(\Delta\mathbf{R}_{\text{ref},i}^{T}\Delta\mathbf{R}_{\text{est},i})-1}{2}\right)\cdot\frac{180}{\pi}$ (32)
where $\Delta\mathbf{R}_{\text{ref},i}=\mathbf{R}_{\text{ref},i+1}\mathbf{R}_{\text{ref},i}^{T}$ and $\Delta\mathbf{R}_{\text{est},i}=\mathbf{R}_{\text{est},i+1}\mathbf{R}_{\text{est},i}^{T}$. The RMSE over all consecutive frame pairs is:
$\text{RPE-R}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n-1}\text{RPE-R}_{i}^{2}}$ (33)
B.4 Implementation Details
Preprocessing. Images are resized to $299\times 299$ pixels for FID/FVD computation and $224\times 224$ pixels for CLIP-based metrics, with appropriate normalization applied. For camera trajectory evaluation, images are resized to $720\times 480$ pixels and uniformly sampled to 20 frames while preserving the first and last frames. Videos are uniformly sampled to 25 frames while preserving the first and last frames. For computational efficiency, FVD calculation further samples to 16 frames, and $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$ processing is limited to 20 frames for long videos.
Evaluation Protocol. For static scenes with multiple reference images, we directly construct the real distribution using all available images. For single-image scenes, we apply minimal augmentation strategies to avoid singular covariance matrices. Dynamic scenes maintain frame correspondence between generated and reference videos to ensure fair comparison. For camera trajectory evaluation, we employ Structure-from-Motion (SfM) to reconstruct camera poses from image sequences, then apply Sim3 alignment to handle scale ambiguity inherent in monocular reconstruction. Trajectory comparisons are performed using the evo toolkit with appropriate alignment and scale correction parameters.

### v2

@@
 Abstract
-Recent video diffusion models demonstrate strong potential in spatial intelligence tasks due to their rich latent world priors.
-However, this potential is hindered by their limited controllability and geometric inconsistency, creating a gap between their strong priors and their practical use in 3D/4D tasks.
-As a result, current approaches often rely on retraining or fine-tuning, which risks degrading pretrained knowledge and incurs high computational costs.
+Recent video diffusion models show immense potential for spatial intelligence tasks due to their rich world priors, but this is undermined by limited controllability, poor spatial-temporal consistency, and entangled scene-camera dynamics.
+Existing solutions, such as model fine-tuning and warping-based repainting, struggle with scalability, generalization, and robustness against artifacts.
 To address this, we propose WorldForge, a training-free, inference-time framework composed of three tightly coupled modules.
-Intra-Step Recursive Refinement introduces a recursive refinement mechanism during inference, which repeatedly optimizes network predictions within each denoising step to enable precise trajectory injection.
-Flow-Gated Latent Fusion leverages optical flow similarity to decouple motion from appearance in the latent space and selectively inject trajectory guidance into motion-related channels.
-Dual-Path Self-Corrective Guidance compares guided and unguided denoising paths to adaptively correct trajectory drift caused by noisy or misaligned structural signals.
+1) Intra-Step Recursive Refinement injects fine-grained trajectory guidance at denoising steps through a recursive correction loop, ensuring motion remains aligned with the target path.
+2) Flow-Gated Latent Fusion leverages optical flow similarity to decouple motion from appearance in the latent space and selectively inject trajectory guidance into motion-related channels.
+3) Dual-Path Self-Corrective Guidance compares guided and unguided denoising paths to adaptively correct trajectory drift caused by noisy or misaligned structural signals.
 Together, these components inject fine-grained, trajectory-aligned guidance without training, achieving both accurate motion control and photorealistic content generation.
-Extensive experiments across diverse benchmarks validate our method’s superiority in realism, trajectory consistency, and visual fidelity.
-This work introduces a novel plug-and-play paradigm for controllable video synthesis, offering a new perspective on leveraging generative priors for spatial intelligence.
-Figure 1: We present WorldForge, a fully training-free framework leveraging a pre-trained video diffusion model for various 3D/4D tasks, such as monocular 3D scene generation (up) and dynamic 4D scene re-rendering (down), enabling precise camera trajectory control and high-quality outputs.
+Our framework is plug-and-play and model-agnostic, enabling broad applicability across various 3D/4D tasks.
+Extensive experiments demonstrate that our method achieves state-of-the-art performance in trajectory adherence, geometric consistency, and perceptual quality, outperforming both training-intensive and inference-only baselines.
+Figure 1: We present WorldForge, a fully training-free framework leveraging a pre-trained video diffusion model for various 3D/4D tasks, such as monocular 3D scene generation (left) and dynamic 4D scene re-rendering (right), enabling precise camera trajectory control and high-quality outputs.
 1 Introduction
-Recent advances in generative modeling, particularly video diffusion models (VDM) (Blattmann et al., 2023; Wan et al., 2025; Yang et al., 2024; Google DeepMind, 2025), have greatly expanded the scope and capabilities of spatial intelligence (Cao et al., 2025) tasks such as 3D and 4D scene understanding (Bahmani et al., 2025a; b), reconstruction (Wang et al., 2025a; Wu et al., 2025; Shi et al., 2024), and generation (Yu et al., 2024c; 2025).
-Trained on massive, diverse video datasets, these models inherently encode rich spatiotemporal priors that capture structural, temporal, and motion-related patterns.
-Leveraging such priors offers significant advantages for achieving realistic and coherent spatial transformations, enabling applications in novel view synthesis (You et al., 2025; Xiao et al., 2025), panoramic video generation (Wang et al., 2024b; Ma et al., 2024a), 3D scene generation (Liu et al., 2024; Liang et al., 2025; Wang et al., 2024c), and dynamic scene reconstruction (Bai et al., 2025a; Yu et al., 2025; Van Hoorick et al., 2024).
-Moreover, VDMs are increasingly used to construct “world models” (Bar et al., 2025; Duan et al., 2025; Bruce et al., 2024), understood as structured internal representations of physical environments that support predictive reasoning, planning, and decision-making in embodied AI systems.
-Despite their strong generative capabilities, current video diffusion models face fundamental limitations, including limited controllability, spatial–temporal consistency, and geometric fidelity, particularly when applied to 3D or 4D tasks (Wang et al., 2024c; He et al., 2024; Ling et al., 2024; Xing et al., 2024).
-While they can be loosely guided with text prompts or reference frames (Kong et al., 2024; Wan et al., 2025), they generally lack the ability to follow precise user-defined motion constraints, such as a specified 6-DoF camera trajectory or object pose evolution (Hu, 2024; Ma et al., 2024b).
-This lack of trajectory-level control is especially problematic in tasks requiring spatial consistency across views, such as novel view synthesis or free-viewpoint rendering.
-Furthermore, these models often entangle scene and camera motion, making it difficult to separate object dynamics from viewpoint changes (Yu et al., 2024c; Liu et al., 2024).
-Consequently, attempts to synthesize videos with fixed scenes or smooth camera paths often result in unintended object deformations or scene instability.
-These limitations hinder their applicability in domains requiring structured spatial reasoning or controllable video generation.
-To handle these limitations, prior works (Jeong et al., 2025; Ren et al., 2025; Yu et al., 2025; Zhang et al., 2025) have explored two main directions.
-The first involves training or fine-tuning the generative backbone on multiview or motion-conditioned data, often with explicit modules to encode trajectory information (Bai et al., 2025a; Xiao et al., 2024; Bai et al., 2025b).
-While this approach can improve alignment and control, it is computationally costly, may generalize poorly to diverse scenes, and risks degrading the model’s pretrained priors during fine-tuning.
-The second line of work adopts a “warping-and-repainting” strategy (Ma et al., 2025b; Liu et al., 2025; Ma et al., 2025a; You et al., 2025), in which input frames are lifted into a partial 3D representation (e.g., via depth estimation (Piccinelli et al., 2024; Yin et al., 2023)), re-projected along a user-defined camera path, and then refined by a generative model to fill missing regions.
-Although more flexible, such methods face notable robustness limitations, as pretrained models are not designed to process out-of-distribution (OOD) (Yu et al., 2024a) inputs such as warped or disoccluded images.
-As a result, they often produce artifacts and missing structures (e.g., suspended components or fragmented geometry).
-Moreover, their bias toward dynamic training data leads to hallucinated motion even in static scenes, undermining view consistency.
-In summary, the combined effects of OOD inputs and dynamic-data bias make it challenging to balance fine-grained controllability with generation quality and generalization, which remains an open problem.
+Recent video diffusion models (VDMs) (Blattmann et al., 2023; Wan et al., 2025; Yang et al., 2024; Google DeepMind, 2025) have significantly advanced spatial intelligence (Cao et al., 2025) tasks like 3D/4D understanding (Bahmani et al., 2025a; b), reconstruction (Wang et al., 2025a; Wu et al., 2025; Shi et al., 2024), and generation (Yu et al., 2024c; 2025).
+Trained on vast video datasets, these models encode rich spatiotemporal priors, enabling realistic spatial transformations for applications like novel view synthesis (You et al., 2025; Xiao et al., 2025), panoramic video (Wang et al., 2024b; Ma et al., 2024a), and dynamic scene reconstruction (Bai et al., 2025a; Yu et al., 2025; Van Hoorick et al., 2024).
+Furthermore, VDMs are increasingly used to build “world models” (Bar et al., 2025; Duan et al., 2025; Bruce et al., 2024), which are structured internal representations that support predictive reasoning in embodied AI.
+Despite their strong priors, VDMs face fundamental limitations, including limited controllability, spatial–temporal consistency, and geometric fidelity, particularly when applied to 3D or 4D tasks (Wang et al., 2024c; He et al., 2024; Ling et al., 2024; Xing et al., 2024).
+They struggle to follow precise motion constraints, such as a 6-DoF camera trajectory (Hu, 2024; Ma et al., 2024b), which undermines spatial consistency in tasks such as novel view synthesis and trajectory control.
+These models also entangle scene and camera motion, causing unintended object deformations and scene instability when viewpoint changes are desired (Yu et al., 2024c; Liu et al., 2024).
+These limitations hinder their use in applications requiring structured spatial reasoning or controllable generation.
+To handle these issues, prior works (Jeong et al., 2025; Ren et al., 2025; Yu et al., 2025; Zhang et al., 2025) have pursued two main directions.
+The first, fine-tuning on motion-conditioned data (Bai et al., 2025a; Xiao et al., 2024; Bai et al., 2025b), can improve control but is computationally costly, generalizes poorly, and risks degrading pretrained priors.
+The second, a ”warping-and-repainting” strategy (Ma et al., 2025b; Liu et al., 2025; Ma et al., 2025a; You et al., 2025), re-projects frames along a new camera path and uses a generative model to fill occlusions.
+Although this approach is more flexible, it lacks robustness because pretrained models handle warped, out-of-distribution (OOD) (Yu et al., 2024a) inputs poorly, often producing artifacts and fragmented geometry.
+Moreover, a bias toward dynamic training data can cause hallucinated motion in static scenes.
+Consequently, balancing fine-grained controllability with generation quality and generalization remains a challenging open problem.
 To address this challenge, we aim to inject precise control into VDMs while preserving their valuable priors.
-For this purpose, we propose a general inference-time guidance paradigm that leverages the rich priors of large-scale VDMs (Blattmann et al., 2023; Wan et al., 2025) in spatial intelligence tasks, such as geometry-aware 3D scene generation and video trajectory control.
-Our method adopts a warping-and-repainting pipeline, in which input frames are warped along a reference trajectory and then used as conditional inputs in the repainting stage.
+For this purpose, we propose a general inference-time guidance paradigm that leverages the rich priors of VDMs in spatial intelligence tasks, such as geometry-aware 3D scene generation and video trajectory control.
+Our method uses a warping-and-repainting pipeline, in which input frames are warped along a reference trajectory and then used as conditional inputs in the repainting stage.
 Building on this, we develop a unified, training-free framework composed of three complementary mechanisms, each designed to address a specific challenge in trajectory-controlled generation.
 First, to ensure the generated motion follows the target trajectory derived from depth-based rendering (Wang et al., 2025b; Piccinelli et al., 2024), we introduce Intra-Step Recursive Refinement (IRR).
-It embeds a micro-scale predict–correct loop within each denoising step: before the next timestep, predicted content in observed regions is replaced with the corresponding ground-truth observations.
-This incremental correction allows trajectory control signals to be injected at every step, enabling fine-grained, stepwise guidance that keeps the motion aligned with the target trajectory.
+It embeds a micro-scale predict–correct loop within each denoising step: before the next timestep, predicted content in observed regions is replaced with the corresponding ground-truth (GT) observations.
+This incremental correction allows trajectory control signals to be injected at every step, enabling fine-grained aligned with the target trajectory.
 Second, we observe that different channels of the VAE-encoded (Kingma & Welling, 2013; Foti et al., 2022) latent representation encode different information, with some channels specializing in appearance and others in motion.
-Directly overwriting all channels when injecting trajectory signals can inadvertently degrade fine-grained visual details.
-To address this, we propose Flow-Gated Latent Fusion (FLF), which leverages optical-flow similarity to selectively and dynamically inject trajectory information only into motion-relevant channels, while leaving appearance-relevant channels unmodified.
-This selective modulation effectively decouples appearance from motion, allowing for precise viewpoint manipulation while preserving content fidelity.
-Finally, while warping-based rendering effectively enforces user-defined trajectories in view synthesis, it inevitably introduces noise stemming from imperfect depth estimation, occlusions, and scene misalignments.
-These imperfections often lead to visual artifacts such as ghosting, structural distortions, and degraded temporal coherence (You et al., 2025).
-To better balance trajectory adherence and generation fidelity, we propose a Dual-Path Self-Corrective Guidance (DSG) strategy.
-Inspired by CFG (Ho & Salimans, 2021), DSG introduces two concurrent denoising pathways during inference: a non-guided path that relies on the model’s internal priors and yields high-fidelity but uncontrolled outputs, and a guided path conditioned on the warped input trajectory, which enforces camera motion but is prone to artifacts.
-By utilizing the difference between these two paths at each denoising step, DSG computes a dynamic correction term that softly adjusts the guided path toward the perceptual quality of the non-guided path.
-This self-corrective mechanism effectively mitigates trajectory-induced degradation while maintaining alignment with the target camera path, thereby improving both the structural integrity and visual quality of the generated video.
-Together, these three mechanisms form a cohesive inference-time guidance framework that achieves robust and precise trajectory control while preserving the generalization ability of VDM priors.
-Our method is fully training-free and plug-and-play, enabling broad applicability across tasks without model retraining.
-It is also model-agnostic and readily adapts to backbones such as Wan 2.1 (Wan et al., 2025) and SVD (Blattmann et al., 2023), underscoring its general utility.
-To validate the effectiveness of the proposed method, we conduct comprehensive experiments on multiple tasks and benchmarks.
-Results demonstrate consistent improvements in trajectory adherence, geometric consistency, and perceptual quality compared to state-of-the-art (SOTA) baselines.
-Our main contributions are summarized as follows:
-  • We introduce a novel, training-free paradigm for leveraging video generative priors in spatial intelligence tasks, enabling precise and stable 3D/4D trajectory control without retraining or fine-tuning.
-  • We design a synergistic inference-time guidance framework integrating Intra-Step Recursive Refinement (IRR) and Flow-Gated Latent Fusion (FLF), achieving accurate trajectory adherence while disentangling motion from content.
-  • We propose Dual-Path Self-Corrective Guidance (DSG), a self-referential correction mechanism that enhances spatial alignment and perceptual fidelity without auxiliary networks or retraining.
-  • We demonstrate, through extensive experiments on diverse datasets and tasks, that our approach achieves state-of-the-art controllability and visual quality, even compared to training-intensive pipelines.
+Directly overwriting all channels when injecting trajectory signals can inadvertently degrade visual details.
+To address this, we propose Flow-Gated Latent Fusion (FLF), which leverages optical-flow similarity to selectively inject trajectory information into channels highly correlated with motion, while leaving appearance-relevant channels unmodified.
+This process effectively decouples appearance from motion, allowing for precise viewpoint manipulation while preserving content fidelity.
+Finally, while warping-based rendering effectively enforces user-defined trajectories, it inevitably introduces noise and visual artifacts stemming from imperfect depth, occlusions, or misalignments (You et al., 2025).
+To balance control with quality, we propose Dual-Path Self-Corrective Guidance (DSG).
+Inspired by CFG (Ho & Salimans, 2021), DSG uses two parallel denoising paths during inference: A non-guided path that relies on the model’s priors to produce high-fidelity but uncontrolled results, and a guided path that follows the warped trajectory, ensuring camera control but risking artifacts.
+At each step, DSG computes the difference between these paths to create a dynamic correction term.
+This term adjusts the guided path toward the higher perceptual quality of the non-guided path.
+This self-corrective mechanism mitigates artifacts from the warped trajectory while maintaining camera control, improving the video’s overall structure and visual quality.
+Together, these three mechanisms form a cohesive inference-time guidance framework for robust and precise trajectory control while preserving VDM priors.
+Our method is training-free and plug-and-play, enabling broad applicability across tasks without model retraining.
+It is also model-agnostic and readily adapts to backbones such as Wan 2.1 (Wan et al., 2025) and SVD (Blattmann et al., 2023).
+Comprehensive experiments on multiple tasks and benchmarks confirm that our approach achieves state-of-the-art (SOTA) results, improving trajectory adherence, geometric consistency, and perceptual quality over leading baselines.
+Our main contributions are:
+  • A novel, training-free paradigm for leveraging VDM priors in spatial intelligence tasks, enabling precise and stable 3D/4D trajectory control without retraining or fine-tuning.
+  • A synergistic inference-time guidance framework integrating Intra-Step Recursive Refinement (IRR) and Flow-Gated Latent Fusion (FLF), achieving accurate trajectory adherence while disentangling motion from content.
+  • Dual-Path Self-Corrective Guidance (DSG), a self-referential correction mechanism that enhances spatial alignment and perceptual fidelity without auxiliary networks or retraining.
+  • Extensive experiments on diverse datasets and tasks show our approach achieves SOTA controllability and visual quality, even compared to training-intensive pipelines.
 Figure 2: Overview of our porposed method.
@@
 2 Related Works
-We review prior work in three areas most relevant to ours: 3D static scene generation, 4D trajectory-controlled video generation, and guidance strategies for generative models.
+We review prior work in three relevant areas: 3D static scene generation, 4D trajectory-controlled video generation, and guidance strategies for generative models.
 3D Static Scene Generation.
-Recent advances in 3D reconstruction (Mildenhall et al., 2020; Kerbl et al., 2023; Song et al., 2024; Gao et al., 2024; Yu et al., 2024d; Müller et al., 2022; Yao et al., 2018) and object-level generation (Poole et al., 2023; Wei et al., 2024; Xiang et al., 2025; Kwak et al., 2024) have achieved strong results, but they lack scene-level priors and struggle with geometric consistency, limiting their scalability.
-VDM (Blattmann et al., 2023; Wan et al., 2025; Kong et al., 2024) offer richer priors and have thus become a foundation for scene generation.
-Approaches vary: some decode 3D scenes from a single image via latent traversal (e.g., Wonderland (Liang et al., 2025)), others fine-tune VDMs with depth-based warped inputs, (e.g., ViewCrafter(Yu et al., 2024c) and See3D (Ma et al., 2025a)), and some methods such as MotionCtrl (Wang et al., 2024c) and TrajectoryAttention (Xiao et al., 2025) embed camera parameters to guide view synthesis .
-More efficient, training-free strategies such as NVS-Solver (You et al., 2025) and ViewExtrapolator (Liu et al., 2024) warp input views and modulate frozen models during sampling.
-Fine-tuning offers controllability but risks eroding pretrained priors and incurs high cost, while training-free methods retain priors and efficiency but must address geometric coherence.
-Our work follows the training-free direction, aiming to ensure both view consistency and controllability.
+While 3D reconstruction (Mildenhall et al., 2020; Kerbl et al., 2023; Song et al., 2024; Gao et al., 2024; Yu et al., 2024d; Müller et al., 2022; Yao et al., 2018) and object generation (Poole et al., 2023; Wei et al., 2024; Xiang et al., 2025; Kwak et al., 2024) are advanced, they often lack scene-level priors.
+VDM (Blattmann et al., 2023; Wan et al., 2025; Kong et al., 2024) provide these priors and are leveraged by decoding scenes from images (Liang et al., 2025), fine-tuning on warped inputs (Yu et al., 2024c; Ma et al., 2025a), or embedding camera parameters (Wang et al., 2024c; Xiao et al., 2025).
+Unlike costly fine-tuning which can corrupt priors, training-free strategies (You et al., 2025; Liu et al., 2024) preserve them but must ensure geometric coherence.
+Our work takes this training-free approach to enhance both consistency and control.
 Trajectory-Controlled Dynamic Video Generation.
-Video synthesis with controllable camera motion generally follows two main paradigms.
-The first is fine-tuning-based (Ma et al., 2025b; Mou et al., 2024; Yu et al., 2024b; Wang et al., 2024c) , where lightweight adapters $a_{\phi}$ (e.g., LoRA (Hu et al., 2022), ControlNet (Zhang et al., 2023)) are trained on video–trajectory pairs to optimize the standard diffusion (Ho et al., 2020) loss.
-Examples include ReCamMaster (Bai et al., 2025a), which retrains on 136K annotated videos, TrajectoryCrafter (Yu et al., 2025) with dual-stream conditioning on source videos and 3D point-cloud renders, GCD (Van Hoorick et al., 2024) using synthetic multi-view videos, and DaS (Gu et al., 2025) incorporating 3D tracking signals into a pre-trained diffusion model for multi-type motion control.
-The second paradigm is warp-and-repaint, which projects source frames with depth to target poses and then repaints the occluded regions (Ma et al., 2025b; Liu et al., 2025; Huang et al., 2025; Tian et al., 2025).
-While flexible, this approach remains vulnerable to noisy warps that cause flicker or distortion—as in training-free baselines such as NVS-Solver (You et al., 2025), which modulates a frozen video diffusion model using warped views at inference time.
-In contrast, our method applies a more powerful inference-time guidance mechanism that extracts trajectory cues and directly steers the diffusion process, achieving precise motion-consistency control without any training.
+One paradigm for controllable video synthesis is to fine-tune lightweight adapters (Ma et al., 2025b; Mou et al., 2024; Yu et al., 2024b; Wang et al., 2024c) like LoRA (Hu et al., 2022) or ControlNet (Zhang et al., 2023) on video-trajectory data, conditioning on diverse inputs (Bai et al., 2025a; Yu et al., 2025; Van Hoorick et al., 2024; Gu et al., 2025).
+Another is the warp-and-repaint strategy (Ma et al., 2025b; Liu et al., 2025; Huang et al., 2025; Tian et al., 2025), which projects and inpaints frames but is prone to artifacts from noisy warps (You et al., 2025).
+Our work uses inference-time guidance to directly steer the diffusion process for precise motion control.
 Guidance and Control for Generative Models.
-A central challenge in diffusion models is how to steer the generative process toward desired outputs.
-Guidance strategies address this by modifying the sampling trajectory to better satisfy conditioning signals.
-The most common is Classifier-Free Guidance (CFG) (Ho & Salimans, 2021), which biases generation toward a target condition by blending conditional and unconditional score predictions.
-While effective, high guidance weights can cause distortions.
-More advanced methods, such as Auto-Guidance and STG (Karras et al., 2024; Hyung et al., 2025; Xu et al., 2023), use an auxiliary model to anticipate and avoid failure modes, whereas Z-sampling (Bai et al., 2025c) alternates denoising and inversion to refine results mid-generation.
-For 3D and 4D synthesis, enforcing viewpoint consistency often follows the fine-tuning and warp-and-repaint strategies discussed above.
-However, the latter remains sensitive to noisy warps (Cai et al., 2024; Wang et al., 2024a), leading to flicker and geometric distortions.
-To this end, we propose Dual-Path self-corrective guidance, which derives a correction signal from the difference between guided and unguided predictions at each step, enhancing trajectory adherence and stability without retraining or per-scene tuning.
+Guidance strategies steer diffusion models toward desired outputs.
+While Classifier-Free Guidance (CFG) (Ho & Salimans, 2021) is common, high weights can cause artifacts.
+More advanced techniques use auxiliary models (Karras et al., 2024; Hyung et al., 2025; Xu et al., 2023) or iterative refinement (Bai et al., 2025c) to improve sampling.
+In 3D/4D synthesis, guidance is used to enforce viewpoint consistency, but warp-based methods often suffer from noise-induced artifacts (Cai et al., 2024; Wang et al., 2024a).
+To address this, we propose DSG.
+It derives a correction signal from the difference between guided and unguided predictions at each step, enhancing trajectory adherence and stability without retraining.
 3 Proposed Methods
-We address the challenge of balancing controllability, visual fidelity, and generalization when applying video diffusion models (VDMs) to 3D/4D tasks.
-Our solution is a training-free framework for trajectory-controlled video generation.
-At its core is an inference-time guidance strategy that steers a pre-trained diffusion model along user-defined trajectories while preserving its intrinsic priors and generative quality.
+We propose an inference-time guidance strategy to balance controllability with visual fidelity for VDMs in 3D/4D tasks.
+Our method is a training-free framework that steers a pretrained model along a user-defined trajectory while preserving its generative priors.
 As shown in Fig.
-2, the framework integrates three complementary components.
-Intra-Step Recursive Refinement (IRR) injects trajectory guidance from observed regions at each denoising step, ensuring consistent control throughout sampling (see Sec.
+2, our framework has three key components.
+First, Intra-Step Recursive Refinement (IRR) injects trajectory guidance from observed regions at each denoising step for consistent control (Sec.
 3.2).
-Flow-Gated Latent Fusion (FLF) refines trajectory injection by decoupling motion and appearance features in the latent space, which prevents content drift and preserves fidelity (see Sec.
+Second, Flow-Gated Latent Fusion (FLF) decouples motion from appearance in the latent space to prevent content drift and preserve fidelity (Sec.
 3.3).
-Dual-Path Self-Corrective Guidance (DSG) enhances stability by comparing guided and unguided predictions, using their difference as a corrective signal to suppress artifacts from noisy priors (see Sec.
+Finally, Dual-Path Self-Corrective Guidance (DSG) uses the difference between guided and unguided paths as a corrective signal to suppress artifacts and improve stability (Sec.
 3.4).
-Together, these modules unlock the model’s latent 3D/4D awareness and enable fine-grained trajectory control without retraining.
+Together, these modules enable fine-grained trajectory control and unlock the model’s latent 3D/4D awareness without any retraining.
 3.1 Preliminaries
@@
 Diffusion Solvers.
-Modern generative modeling is dominated by two paradigms: diffusion models (Ho et al., 2020; Song et al., 2020a) and flow-based models (Lipman et al., 2022).
-Under the SDE view, diffusion models admit a deterministic ODE limit that connects them to flow-based formulations through reparameterization (Gao et al., 2025).
-The detailed derivation of this equivalence is provided in Appendix A.
-In this section, we take diffusion models as representative examples.
-Consider the widely used DDIM sampler (Song et al., 2020a): it recovers the clean sample $\mathbf{x}_{0}$ by reversing the forward noising of a Gaussian prior $\mathbf{x}_{T}$.
+Generative models are largely diffusion (Ho et al., 2020) or flow-based (Lipman et al., 2022).
+Under the SDE view, diffusion models have a deterministic ODE limit that connects to flow models via reparameterization (Gao et al., 2025) (The detailed derivation is provided in Appendix A).
+We use the popular DDIM sampler (Song et al., 2020a) as an example to illustrate the sampling process: it recovers the clean sample $\mathbf{x}_{0}$ by reversing the forward noising of a Gaussian prior $\mathbf{x}_{T}$.
 Given a noise-prediction network $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$, the sampler estimates an intermediate signal $\hat{\mathbf{x}}_{0}$ from the current state $\mathbf{x}_{t}$:
@@
 $\mathbf{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\,\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},t)+\sqrt{1-\bar{\alpha}_{t-1}}\,\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t).$ (2)
-Iterating this update from $t=T$ to $t=0$ produces the final sample $\mathbf{x}_{0}$.
+Iterating from $t=T$ to $t=0$ yields the final sample $\mathbf{x}_{0}$.
 Our method intervenes at this stage by modifying $\hat{\mathbf{x}}_{0}$ to enforce trajectory control.
-Notably, other popular solvers, such as UniPC, EDM, and PNDM (Zhao et al., 2023; Karras et al., 2022; Liu et al., 2022a), also compute $\hat{\mathbf{x}}_{0}$ directly or can recover it via a parameterized transformation, so our framework is broadly compatible.
-Classifier Free Guidance. To improve fidelity to the condition, CFG (Ho & Salimans, 2021) adjusts the score function during sampling:
-$\tilde{\bm{\epsilon}}_{\theta}(\mathbf{x}_{t},t)=\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,\phi)+\omega_{\text{CFG}}\cdot\left[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,\phi)\right],$ (3)
-where $\omega_{\text{CFG}}$ is the guidance weight, with $\mathbf{c}$ and $\phi$ denoting the conditional and unconditional branches, respectively.
+Notably, other popular solvers, such as UniPC (Zhao et al., 2023), EDM (Karras et al., 2022), and PNDM (Liu et al., 2022a), also compute $\hat{\mathbf{x}}_{0}$ directly or can recover it via a parameterized transformation, so our framework is broadly compatible.
+Since our experiments primarily use the flow-based Wan2.1 (Wan et al., 2025) model, we will subsequently detail our algorithm using a flow-based formulation.
+Classifier Free Guidance. To improve fidelity to the condition, CFG (Ho & Salimans, 2021) adjusts the network prediction during sampling as:
+$\tilde{\bm{\epsilon}}_{\theta}(\mathbf{x}_{t},t)=\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})+\omega_{\text{CFG}}\cdot\left[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,\phi)\right],$ (3)
+where $\omega_{\text{CFG}}$ is the guidance weight, with $\mathbf{c}$ and $\phi$ denoting the conditional and unconditional inputs, respectively.
 This interpolates conditional and unconditional scores to steer the sampling trajectory.
@@
 3.1.2 Trajectory Control via Depth-Based Warping
-Our framework uses a warping-and-repainting strategy, embedding geometric cues through depth-based view warping.
-For reliable depth estimation to support warping into new views, we employ depth prediction networks that takes one or more input images $\{\mathbf{I}_{i}\}_{i=1}^{N}$ and estimates corresponding camera poses and depth maps:
-$f:\{\mathbf{I}_{i}\}_{i=1}^{N}\rightarrow\{\mathbf{P}_{i},\mathbf{D}_{i}\},$ (4)
-where the depth maps $\mathbf{D}_{i}$ and poses $\mathbf{P}_{i}$ are then used to warp the source views to target poses.
-Formally, the warping function $\mathcal{W}$ projects source frames $\mathbf{I}_{src}$ with depth $\mathbf{D}_{src}$ from pose $\mathbf{P}_{src}$ to target pose $\mathbf{P}_{tar}$, producing partial target views $\mathbf{I}^{\prime}_{tar}$ and validity masks $\mathbf{M}_{tar}$, which indicate the visible pixels in warped views:
+Our framework controls trajectories using a depth-based warping-and-repainting strategy.
+First, a depth-prediction network estimates camera poses and depth maps $(\mathbf{P}_{q},\mathbf{D}_{q})$ from single image $\mathbf{I}$} or image sequence $\{\mathbf{I}_{q}\}_{q=1}^{N}$ via a function $f:\{\mathbf{I}_{q}\}_{q=1}^{N}\rightarrow\{\mathbf{P}_{q},\mathbf{D}_{q}\}$.
+Next, a warping operator $\mathcal{W}$ uses these estimates to project a source frame $\mathbf{I}_{src}$ with depth $\mathbf{D}_{src}$ from pose $\mathbf{P}_{src}$ to a target pose $\mathbf{P}_{tar}$.
+This yields a partial target view $\mathbf{I}^{\prime}_{tar}$ and a validity mask $\mathbf{M}_{tar}$ indicating visible pixels:
 $(\mathbf{I}^{\prime}_{tar},\mathbf{M}_{tar})=\mathcal{W}(\mathbf{I}_{src},\mathbf{D}_{src},\mathbf{P}_{src},\mathbf{P}_{tar}).$ (4)
-The warped frames $\mathbf{I}^{\prime}_{tar}$ and masks $\mathbf{M}_{tar}$ provide trajectory-aware observations that serve as the basis for generating videos along arbitrary target poses $\mathbf{P}_{tar}$, though limited to regions visible in source views.
-With these preliminaries, we now present our method.
-Our goal is to use the trajectory control introduced in this section to guide video generation in VDM.
-To achieve precise and consistent control, we design three modules: Intra-Step Recursive Refinement, Flow-Gated Latent Fusion, and Dual-Path Self-Corrective Guidance.
-Together, they complement each other to ensure accurate trajectory guidance while preserving high-quality synthesis.
+The resulting warped frames and masks guide the VDMs along the target poses $\mathbf{P}_{tar}$.
+This guidance is limited to regions visible in the source views.
+With these preliminaries, we use this trajectory control to guide video generation in VDMs.
 3.2 Intra-Step Recursive Refinement
@@
 As noted in Sec.
-3.1.1, the denoising process produces an intermediate variable $\hat{\mathbf{x}}_{0}^{(i)}$, a coarse estimate of the final output and the baseline for later steps.
-Building on this, IRR modifies $\hat{\mathbf{x}}_{0}^{(i)}$ to impose trajectory constraints, ensuring that generation follows the desired path.
-IRR operates within the update process of Eq.
+3.1.1, the denoising process produces an intermediate variable $\hat{\mathbf{x}}_{0}^{(t)}$, a coarse estimate of the final output and the baseline for later steps,where $t$ denotes the current timestep.
+IRR modifies $\hat{\mathbf{x}}_{0}^{(t)}$ to impose trajectory constraints, ensuring that generation follows the desired path.
+IRR operates within the updates of Eq.
 (1) and Eq.
 (2).
-Given the one-step denoised sample $\hat{\mathbf{x}}_{0}^{(i)}$ from Eq.
-(1), we fuse it with the trajectory latent $\mathbf{Z}_{\text{traj}}$, obtained by encoding the warped frames from Eq.
-(5) into the latent space.
-We then add small Gaussian noise $\bm{\epsilon}$ to obtain a modified latent $\mathbf{x}_{t_{i}}^{\prime}$:
-$\mathbf{x}_{t_{i}}^{\prime}=(1-w(\sigma))\,\mathbf{F}(\hat{\mathbf{x}}_{0}^{(i)},\mathbf{Z}_{\text{traj}})+w(\sigma)\,\cdot\bm{\epsilon},$ (6)
-where $\bm{\epsilon}=\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is the initial Gaussian noise used in sampling.
-To return the fused result $\mathbf{x}_{t_{i}}^{\prime}$ into the denoising process while injecting trajectory information, we reintroduce noise $\bm{\epsilon}$.
-Here $0<w(\cdot)<1$ is a user-defined weighting factor derived from the noise schedule $\sigma$, which controls the strength of the reintroduced noise.
-$\mathbf{F}(\cdot)$ is a mask-based fusion function defined as $\mathbf{F}(\hat{\mathbf{x}}_{0}^{(i)},\mathbf{Z}_{\text{traj}})=\mathbf{M}\cdot\mathbf{Z}_{\text{traj}}+(1-\mathbf{M})\cdot\hat{\mathbf{x}}_{0}^{(i)},$ where $\mathbf{M}$ is the binary validity mask from Eq.
+Given the one-step denoised sample $\hat{\mathbf{x}}_{0}^{(t)}$ from Eq.
+(1), we fuse it with the trajectory latent $\mathbf{Z}_{\text{traj}}$, obtained by encoding the warped frames of Eq.
+(4) into latent space.
+We then add Gaussian noise $\bm{\epsilon}$ to obtain the modified latent $\mathbf{x}_{t}^{\prime}$:
+$\mathbf{x}_{t}^{\prime}=(1-w(\sigma))\,\mathbf{F}(\hat{\mathbf{x}}_{0}^{(t)},\mathbf{Z}_{\text{traj}})+w(\sigma)\,\cdot\bm{\epsilon},$ (5)
+where $\mathbf{F}(\hat{\mathbf{x}}_{0}^{(t)},\mathbf{Z}_{\text{traj}})=\mathbf{M}\cdot\mathbf{Z}_{\text{traj}}+(1-\mathbf{M})\cdot\hat{\mathbf{x}}_{0}^{(t)}$ copies observable warped content from $\mathbf{Z}_{\text{traj}}$ into the corresponding locations of $\hat{\mathbf{x}}_{0}^{(t)}$ using the binary mask $\mathbf{M}$ from Eq.
+(4); and $\bm{\epsilon}=\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is a randomly sampled Gaussian noise used to re-noise the fused latent $\mathbf{F}(\hat{\mathbf{x}}_{0}^{(t)},\mathbf{Z}_{\text{traj}})$ so that it re-enters the denoising schedule with the injected trajectory.
+The re-noising is controlled by a scheduler $w(\sigma)$, and the specific strategy can differ for various diffusion and flow models.
+The re-noised latent $\mathbf{x}_{t}^{\prime}$ is then fed into the network $\bm{\epsilon}_{\theta}$, replacing the original $\mathbf{x}_{t}$ in Eq.
+(1) and Eq.
+(2) for the next sampling step.
+In summary, IRR embeds a micro predictor–corrector at each denoising step.
+By updating $\hat{\mathbf{x}}_{0}^{(t)}$ with explicit trajectory cues, it continually corrects the sampling path and ensures that synthesis follows the target trajectory precisely.
+3.3 Flow-Gated Latent Fusion
+In the IRR process, overwriting all latent channels with trajectory information degrades visual quality because VAE latents encode both motion and appearance.
+The indiscriminate update in Eq.
+(5) injects noise into appearance-focused channels.
+To address this, we propose Flow-Gated Latent Fusion (FLF), a method that identifies and updates latent channels with high motion relevance.
+To select motion-related channels, we use an optical-flow-based scoring scheme since flow directly reflects inter-frame motion.
+First, for each channel $c$ of the latent $\hat{\mathbf{x}}_{0}^{(t)}$ at timestep $t$, we compute the optical flow between consecutive frames to get a predicted flow $\mathcal{F}^{(t,c)}_{\text{pred}}$.
+The resulting map for each channel has a shape of $2,\tau,H,W]$ (flow vectors, frame pairs, spatial dimensions).
+Second, we compute a GT reference flow $\mathcal{F}^{(t,c)}_{\text{gt}}$ from the target trajectory latent $\mathbf{Z}_{\text{traj}}$ in the same frame-by-frame manner.
+Finally, we compare the two flows within the visible regions defined by the mask $\mathbf{M}^{(c)}$.
+The comparison relies on three popular optical flow metrics (Teed & Deng, 2020): Masked End-point Error (M-EPE), which measures the Euclidean distance between predicted and GT flow vectors; Masked Angular Error (M-AE), which measures their angular difference; and Outlier Percentage (Fl-all), which calculates the fraction of unreliable pixels.
+We combine the normalized metrics to calculate a motion similarity score $S^{(t,c)}$ for each channel in each timestep $t$:
+$S^{(t,c)}=\sum_{k\in\{\text{E, A, F}\}}\gamma_{k}\!\left(1-{\text{Norm}_{k}}^{(t,c)}\right),$ (6)
+where $\text{Norm}_{k}^{(t,c)}\!\in[0,1]$ are the normalized errors from M-EPE, M-AE, and Fl-all, and $\gamma_{k}$ are the weighting factors.
+The detailed calculation and normalization methods for these metrics are provided in the Appendix C.
+Higher $S^{(t,c)}$ means better flow alignment and stronger motion evidence.
+To adaptively set the motion similarity threshold for each scene, we select motion-relevant channels using a dynamic threshold $\delta^{(t)}=\mu_{S}^{(t)}-\lambda^{(t)}\sigma_{S}^{(t)}$, where $\mu_{S}^{(t)}$ and $\sigma_{S}^{(t)}$ are the mean and standard deviation of all channel scores at step $t$.
+We schedule $\lambda^{(t)}$ to create a loose-to-tight selection over time.
+This matches the generative process, where early steps define broad structures and later steps handle fine details.
+Consequently, we guide more channels initially for structural integrity and fewer channels toward the end to preserve high-fidelity details.
+Finally, the latent update selectively fuses the trajectory information $\mathbf{Z}_{\text{traj}}$ into the selected high motion-relevance channels:
+$\mathbf{FLF}(\hat{\mathbf{x}}_{0}^{(t)},\mathbf{Z}_{\text{traj}})=\begin{cases}\mathbf{M}^{(c)}\cdot\mathbf{Z}_{\text{traj}}^{(c)}+\left(1-\mathbf{M}^{(c)}\right)\cdot\hat{\mathbf{x}}_{0}^{(t,c)},&\text{if}\;S^{(t,c)}\geq\delta^{(t)}\\ \hat{\mathbf{x}}_{0}^{(t,c)},&\text{otherwise}.\end{cases}$ (7)
+This FLF operator replaces the global update in Eq.
+(5), resulting in a more precise fusion rule: This FLF operator replaces the operator $\mathbf{F}$ in Eq.
+(5), resulting in a more precise fusion rule.
+In summary, FLF provides fine-grained trajectory control while preserving model priors and synthesis quality.
+Unlike methods that restart the sampling schedule (Xu et al., 2023) or globally update the entire latent (Liu et al., 2024), FLF integrates with our IRR framework to apply selective, per-step guidance, effectively decoupling motion and appearance for precise control.
+3.4 Dual-Path Self-Corrective Guidance
+Trajectory latents $\mathbf{Z}_{\text{traj}}$ obtained by warping often contain distortions from depth errors, occlusions, or misalignments, which degrades synthesis quality.
+To mitigate this, we draw inspiration from CFG (Ho & Salimans, 2021).
+Conceptually, in the context of a flow model, CFG treats the unconditional prediction as a “bad” direction $\mathbf{v}_{\text{uncon}}$ and the conditional one as a “good” direction $\mathbf{v}_{\text{con}}$ (Karras et al., 2024).
+It then finds a “better” direction by pushing the “good” one away from the “bad” one.
+Based on this idea, we propose Dual-Path Self-Corrective Guidance (DSG).
+At each iteration, IRR produces two velocity fields.
+The unguided velocity $\mathbf{v}_{t}^{\text{ori}}$ (from the original latent $\mathbf{x}_{t}$) stays on the data manifold with high fidelity but ignores the trajectory, which we consider a “bad” direction for control.
+The guided velocity $\mathbf{v}_{t}^{\text{traj}}$ (from the corrected latent $\mathbf{x}^{\prime}_{t}$) may be noisy but follows the trajectory, which we consider a “good” direction.
+DSG uses the difference between them to find a “better” path.
+However, we emperically find the difference between our $\mathbf{v}_{t}^{\text{traj}}$ and $\mathbf{v}_{t}^{\text{ori}}$ is far greater than that between the $\mathbf{v}_{\text{con}}$ and $\mathbf{v}_{\text{uncon}}$ in standard CFG.
+In extensive tests, we observed that the cosine similarity $\alpha_{t}=(\mathbf{v}_{t}^{\text{traj}}\cdot\mathbf{v}_{t}^{\text{ori}})/(\lVert\mathbf{v}_{t}^{\text{traj}}\rVert\cdot\lVert\mathbf{v}_{t}^{\text{ori}}\rVert)$ between our two paths is typically between 0.3–0.6 (an angle of 50°–70°).
+In contrast, the cosine similarity between $\mathbf{v}_{\text{con}}$ and $\mathbf{v}_{\text{uncon}}$ is nearly 1 (an angle close to 0°), an observation consistent with reports by Wan et al.
 (2025).
-It copies observable warped content from $\mathbf{Z}_{\text{traj}}$ into the corresponding locations of $\hat{\mathbf{x}}_{0}^{(i)}$, while leaving unobserved regions unchanged.
-In summary, IRR embeds a micro predictor–corrector at each denoising step.
-By updating $\hat{\mathbf{x}}_{0}^{(i)}$ with explicit trajectory cues, it continually corrects the sampling path and ensures that synthesis follows the target trajectory precisely.
-3.3 Flow-Gated Latent Fusion
-In the IRR process, directly overwriting all latent channels with trajectory information often degrades visual quality.
-This is because VAE latents tend to encode different types of information: some channels mainly capture appearance, while others relate more to motion.
-Using Eq.
-(6) to replace all channels indiscriminately introduces noise into appearance-dominant ones.
-To avoid this, we introduce Flow-Gated Latent Fusion (FLF), which decouples motion and appearance features in the latent space and selectively injects trajectory guidance into motion-relevant channels.
-To separate motion from appearance and identify motion-relevant channels, we introduce a flow-based scoring scheme.
-Optical flow captures pixel-wise motion between frames and is widely used to describe temporal dynamics.
-By comparing predicted and reference flows, we estimate how strongly each latent channel encodes motion.
-Given the predicted latent $\hat{\mathbf{x}}_{0}^{(i)}$ from IRR and the reference trajectory latent $\mathbf{Z}_{\text{traj}}$, we compute per-channel flows using the Farnebäck algorithm.
-For each channel $c$, we obtain predicted flow $\mathcal{F}^{(c)}_{\text{pred}}(x,y,\tau)=(u^{(c)}_{\text{pred}}(x,y,\tau),v^{(c)}_{\text{pred}}(x,y,\tau))$ and reference ground truth (GT) flow $\mathcal{F}^{(c)}_{\text{gt}}(x,y,\tau)=(u^{(c)}_{\text{gt}}(x,y,\tau),v^{(c)}_{\text{gt}}(x,y,\tau))$, restricted to valid regions $\mathbf{M}^{(c)}(x,y,\tau)$ from Eq.
-(5).
-Here $(x,y)$ are pixel coordinates, $\tau$ is the time index, and $u_{*},v_{*}$ are the horizontal and vertical components.
-To evaluate motion relevance, we design a normalized score that combines three masked metrics: Masked End-point Error (M-EPE), Masked Angular Error (M-AE), and Outlier Percentage (Fl-all).
-M-EPE is the average distance between the vectors of the predicted flow and GT flow; smaller values mean better alignment.
-M-AE measures the angular difference between flows; smaller values mean more accurate directions.
-Fl-all is the percentage of pixels with large deviations; lower values mean fewer outliers and higher reliability.
-Together, these metrics provide a robust estimate of each channel’s motion relevance.
-The formulas are given below:
-M-EPE: Given the set $\Omega$ of all valid pixels $(x,y,\tau)$ where $\mathbf{M}^{(c)}=1$, we compute the per-channel $L_{2}$ distance between predicted and GT flows and normalize by the number of valid pixels $|\Omega|$:
-$\text{M-EPE}_{c}=\frac{1}{|\Omega|}\sum_{(x,y,\tau)\in\Omega}\left\|\mathcal{F}^{(c)}_{\text{pred}}(x,y,\tau)-\mathcal{F}^{(c)}_{\text{gt}}(x,y,\tau)\right\|_{2}.$ (7)
-M-AE: Similar to M-EPE, for each valid pixel in $\Omega$ we compute the cosine similarity between $\mathcal{F}^{(c)}_{\text{pred}}$ and $\mathcal{F}^{(c)}_{\text{gt}}$, and use it to measure angular error:
-$\text{M-AE}_{c}=\frac{1}{|\Omega|}\sum_{(x,y,\tau)\in\Omega}\arccos\left(\frac{\mathcal{F}^{(c)}_{pred}(x,y,\tau)\cdot\mathcal{F}^{(c)}_{gt}(x,y,\tau)}{\|\mathcal{F}^{(c)}_{pred}(x,y,\tau)\|\cdot\|\mathcal{F}^{(c)}_{gt}(x,y,\tau)\|}\right).$ (8)
-Fl-all: A valid pixel in $\Omega$ is an outlier if its M-EPE exceeds $3$ or its relative error exceeds $5\%$.
-Fl-all (denoted as $\text{F}_{c}$) is the percentage of outliers among valid pixels; lower values mean higher reliability.
-To unify the scales of the three metrics and simplify score computation, we normalize $\text{M-EPE}_{c},\text{M-AE}_{c},\text{F}_{c}$ to $[0,1]$, denoted as $\text{E}^{\text{Norm}}_{c},\text{A}^{\text{Norm}}_{c},$ and $\text{F}^{\text{Norm}}_{c}$, respectively:
-$\text{E}^{\text{Norm}}_{c}=\min(\text{M-EPE}_{c}/n_{1},1),\quad\text{A}^{\text{Norm}}_{c}=\min(\text{M-AE}_{c}/n_{2},1),\quad\text{F}^{\text{Norm}}_{c}=\min(\text{F}_{c}/n_{3},1),$ (9)
-where $n_{1},n_{2},n_{3}$ are normalization factors. We then define the final flow score for each channel as $S$, which directly reflects the consistency between the predicted and GT flows:
-$S^{(c)}=\gamma_{1}\!\left(1-\text{E}^{\text{Norm}}_{c}\right)+\gamma_{2}\!\left(1-\text{A}^{\text{Norm}}_{c}\right)+\gamma_{3}\!\left(1-\text{F}^{\text{Norm}}_{c}\right),$ (10)
-where $\gamma_{1}+\gamma_{2}+\gamma_{3}=1$.
-The flow score $S^{(c)}$ measures the motion relevance of channel $c$.
-A higher score means its predicted flow is closer to the ground truth and thus encodes more motion.
-A lower score indicates misalignment and suggests that the channel encodes more non-motion information, such as appearance.
-This score enables us to decouple motion and appearance channels in the latent space.
-To select motion-relevant channels, we set a dynamic threshold:
-$\delta^{(i)}=\mu_{S}^{(i)}-\lambda^{(i)}\sigma_{S}^{(i)},$ (11)
-where $\mu_{S}^{(i)}$ and $\sigma_{S}^{(i)}$ are the mean and standard deviation of $\{S^{(i,c)}\}$ at step $i$.
-Channels with scores above $\delta^{(i)}$ are regarded as motion-relevant and only these channels receive trajectory guidance.
-The selection sets are:
-$\mathcal{C}_{\text{sel}}^{(i)}=\{\,c\mid S^{(i,c)}\geq\delta^{(i)}\,\}.$
-Only these channels are updated with trajectory latents $\mathbf{Z}_{\text{traj}}$.
-The parameter $0<\lambda^{(i)}<1$ controls the strictness of selection and decreases during denoising.
-Early steps therefore replace more channels for stronger trajectory adherence, while later steps replace fewer channels to preserve appearance details.
-Finally, the latent update rule for $\hat{\mathbf{x}}_{0}^{(i)}$ is:
-$\mathbf{FLF}(\hat{\mathbf{x}}_{0}^{(i)},\mathbf{Z}_{\text{traj}})^{(c)}=\begin{cases}\mathbf{M}^{(c)}\cdot\mathbf{Z}_{\text{traj}}^{(c)}+\left(1-\mathbf{M}^{(c)}\right)\cdot\hat{\mathbf{x}}_{0}^{(i,c)},&c\in\mathcal{C}_{\text{sel}}^{(i)},\\[2.0pt] \hat{\mathbf{x}}_{0}^{(i,c)},&\text{otherwise}.\end{cases}$ (12)
-where $\mathbf{FLF}(\cdot)$ denotes the flow-score–based selective fusion.
-We use it to replace $\mathbf{F}(\cdot)$ in Eq.
-(6), which overwrote all channels without distinction.
-The updated fusion rule is:
-$\mathbf{x}_{t_{i}}^{\prime}=(1-w(\sigma))\,\mathbf{FLF}(\hat{\mathbf{x}}_{0}^{(i)},\mathbf{Z}_{\text{traj}})+w(\sigma)\,\cdot\bm{\epsilon},$ (13)
-In summary, FLF enables fine-grained trajectory injection while preserving model priors and synthesis quality.
-Unlike Restart Sampling (Xu et al., 2023), which restarts the schedule, or ViewExtrapolator (Liu et al., 2024), which injects trajectory data directly into $\hat{\mathbf{x}}_{0}^{(i)}$, our method uses IRR to apply fine-grained guidance at every timestep.
-Combined with FLF, it separates motion from appearance and delivers precise, high-quality trajectory control at inference.
-3.4 Dual-Path Self-Corrective Guidance
-Trajectory latents $\mathbf{Z}_{\text{traj}}$ obtained by warping along the target motion often suffer from distortions due to depth errors, occlusions, or misalignments.
-Such artifacts are common in warp-based methods and degrade synthesis quality.
-To address this, we draw inspiration from CFG (Ho & Salimans, 2021), which interpolates between conditional and unconditional predictions.
-Instead of using conditioning, we exploit the discrepancy between two denoising paths in IRR to obtain a more reliable direction.
-Based on this idea, we propose Dual-Path Self-Corrective Guidance (DSG).
-At each iteration, IRR produces two velocity fields: the unguided $\mathbf{v}_{t_{i}}^{\text{ori}}$ from the original latent $\mathbf{x}_{t_{i}}$, and the guided $\mathbf{v}_{t_{i}}^{\text{traj}}$ from the corrected latent $\mathbf{x}^{\prime}_{t_{i}}$ obtained by injecting trajectory cues.
-DSG leverages their difference to form a correction signal that steers denoising toward a higher-quality path.
-At each step $t_{i}$, IRR and FLF generate two velocity fields: $\mathbf{v}_{t_{i}}^{\text{ori}}$ and $\mathbf{v}_{t_{i}}^{\text{traj}}$.
-The guided velocity $\mathbf{v}_{t_{i}}^{\text{traj}}$ enforces trajectory adherence but may add noise from imperfect warping, while the unguided velocity $\mathbf{v}_{t_{i}}^{\text{ori}}$ stays on the data manifold with higher fidelity but ignores trajectory cues.
-To exploit these two directions for a better denoising path, we propose DSG, which combines them to balance control and fidelity.
-Unlike CFG (Eq.
-(3)) (Ho & Salimans, 2021), which interpolates unconditional and conditional predictions, DSG uses two trajectory-specific complementary branches designed for motion control.
-The corrected velocity $\mathbf{v}_{t_{i}}^{\text{corr}}$ is then computed as:
-$\mathbf{v}_{t_{i}}^{\text{corr}}=\mathbf{v}_{t_{i}}^{\text{traj}}+\rho\cdot\beta_{t_{i}}\bigl(\mathbf{v}_{t_{i}}^{\text{traj}}-\mu\cdot\alpha_{t_{i}}\cdot\mathbf{v}_{t_{i}}^{\text{ori}}\bigr),$ (14)
-where $\rho$ controls guidance strength, $\mu$ normalizes magnitudes, $\alpha_{t_{i}}=(\mathbf{v}_{t_{i}}^{\text{traj}}\cdot\mathbf{v}_{t_{i}}^{\text{ori}})/(\lVert\mathbf{v}_{t_{i}}^{\text{traj}}\rVert\cdot\lVert\mathbf{v}_{t_{i}}^{\text{ori}}\rVert)$ is the cosine similarity between guided and unguided velocities, and $\beta_{t_{i}}=\sin(\arccos(\alpha_{t_{i}}))$ is the corresponding sine similarity.
-This adaptive weighting amplifies corrections when the two paths diverge (lower $\alpha_{t_{i}}$, higher $\beta_{t_{i}}$), pulling the result toward the guided direction, and reduces corrections when they agree (higher $\alpha_{t_{i}}$, lower $\beta_{t_{i}}$), preserving the model’s natural prediction.
-In score-based (Lipman et al., 2022) terms, $\mathbf{v}_{t_{i}}^{\text{ori}}$ follows the pretrained data manifold, while $\mathbf{v}_{t_{i}}^{\text{traj}}$ may deviate due to injected trajectory signals.
-DSG applies cosine-weighted interpolation between the two, dynamically combining them to retain trajectory guidance while projecting the velocity back toward the manifold.
-This improves generation quality by suppressing off-manifold drift.
-As a result, $\mathbf{v}_{t_{i}}^{\text{corr}}$ drives the sample along the desired motion path while preserving visual fidelity to the model priors.
-Subsequent experiments confirm the effectiveness of this mechanism.
+Therefore, directly applying the CFG formula fails in our case.
+To address this large angular difference, we modify the guidance formula to only use the component of the “good” direction that is orthogonal to the “bad” direction.
+This is achieved by projecting $\mathbf{v}_{t}^{\text{traj}}$ onto $\mathbf{v}_{t}^{\text{ori}}$ (after normalizing them to equal length) and taking the difference, which avoids the adverse effects of their large directional divergence:
+$\mathbf{v}_{t}^{\text{corr}}=\mathbf{v}_{t}^{\text{traj}}+\rho\cdot\beta_{t}\bigl(\mathbf{v}_{t}^{\text{traj}}-\alpha_{t}\cdot\mathbf{v}_{t}^{\text{ori}}\bigr),$ (8)
+where $\rho$ controls guidance strength, $\alpha_{t}$ is the cosine similarity, and $\beta_{t}=\sqrt{1-\alpha_{t}^{2}}$ is its sine counterpart.
+The role of $\beta_{t}$ is to adaptively scale the guidance: it amplifies the correction when the paths diverge (low $\alpha_{t}$, high $\beta_{t}$) and reduces it when they agree, preserving the model’s natural prediction.
+In essence, DSG robustly balances precise trajectory control with high visual fidelity.
+Its adaptive cosine-weighting is crucial for handling the large angular difference between the guided and unguided paths.
+It suppresses artifacts by applying strong corrections when the paths diverge, while preserving the model’s natural predictions when they align.
+This ensures the final corrected velocity, $\mathbf{v}_{t}^{\text{corr}}$, steers the sample along the target motion path while maintaining the quality of the model’s priors.
+The effects and details of DSG are presented in Appendix D.
 4 Experiments
@@
 Subsequently, we demonstrate the performance of our method on 3D scene generation and 4D trajectory control in Sec.
-4.2 and Sec.
-4.3, respectively.
+4.2.
 Finally, we conduct a series of ablation studies in Sec.
 4.3 to validate the effectiveness of each component of our approach.
+More results and details are provided in Appendix E.
+Figure 3: Qualitative comparison for 3D novel view synthesis and 4D trajectory-controlled re-rendering.
+(Left) For 3D scene generation from a single image, our method produces more consistent and plausible content compared to SOTA baselines.
+(Right) For 4D re-rendering, our approach leverages model priors to avoid common artifacts like floating heads and flattened faces, yielding more realistic results.
+Overall, our training-free guidance demonstrates superior performance in both static and dynamic scenes.
+Zoom in for more details.
 4.1 Implementation Details
-Our framework is a training-free, inference-time optimization method that steers pre-trained video diffusion models for precise camera control without additional training or fine-tuning.
-It introduces no significant computational overhead beyond the base model’s inference requirements.
+Our framework is a training-free method that steers pre-trained VDMs for precise camera control, with no additional training or fine-tuning.
+Compared with pre-trained VDMs, our method incurs no training cost, while the inference time increases by approximately 40–50%, mainly due to the IRR module (see Appendix E.1 for detailed efficiency analysis).
 […]
 Our pipeline follows a warp-and-repaint design.
-For warping, we test several depth estimation models, such as VGGT (Wang et al., 2025b), UniDepth (Piccinelli et al., 2024), Mega-SaM (Li et al., 2025), and DepthCrafter (Hu et al., 2025).
-The method adapts well to all, benefiting from the strong world priors of the underlying video model.
+For warping, we test several depth estimation models, including VGGT (Wang et al., 2025b), UniDepth (Piccinelli et al., 2024), Mega-SaM (Li et al., 2025), and DepthCrafter (Hu et al., 2025).
+The method adapts well across these choices, benefiting from the VDM’s strong world priors; see Appendix E.3 for depth-estimator ablations.
 Test Datasets and Metrics.
-For single-view 3D scene generation, we adopt the widely used LLFF (Mildenhall et al., 2019), Tanks and Temples (Knapitsch et al., 2017), and MipNeRF 360 (Barron et al., 2022) datasets, selecting half of the scenes from each for evaluation.
-We also test on diverse internet, real-world, and AI-generated images.
-Perceptual quality is assessed using FID (Heusel et al., 2017) and $\mathrm{CLIP}_{\mathrm{sim}}$ (Radford et al., 2021) similarity.
-For 4D trajectory control, we compare with SOTA methods on their respective benchmarks, which include challenging real-world video clips with varied camera trajectories.
-Performance is measured with FVD (Unterthiner et al., 2018) for temporal quality and $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$ for source–target consistency.
-In addition, we assess camera trajectory accuracy using three standard metrics: Absolute Trajectory Error (ATE), which captures the global alignment between estimated and reference paths; Relative Pose Error – Translation (RPE-T), which reflects short-term consistency of translational motion; and Relative Pose Error – Rotation (RPE-R), which quantifies local orientation accuracy.
-Together, these metrics provide a comprehensive evaluation of both global fidelity and local smoothness of camera trajectories.
-4.2 3D Scene Generation
-Figure 3: Qualitative comparison of novel view synthesis from a single input image.
-The first row shows the input frame and its depth-based warped views, where disoccluded regions appear as holes.
-Compared to existing SOTA methods, our approach produces more consistent scene content under novel viewpoints, with improved image detail, trajectory accuracy, and structural plausibility.
-We compare our method with SOTA approaches for novel view synthesis, which can be grouped into two categories.
-The first includes methods requiring model-specific training or fine-tuning, such as ViewCrafter (Yu et al., 2024c), which trains a VDM on iteratively refined point clouds; TrajectoryCrafter (Yu et al., 2025), which learns motion control via dedicated training; TrajectoryAttention (Xiao et al., 2025), which injects motion through a specialized attention module; and See3D(Ma et al., 2025a), a video-based model for scene reconstruction.
-The second category comprises training-free methods, such as NVS-Solver (You et al., 2025), which modulates a pre-trained model using warped views, and ViewExtrapolator (Liu et al., 2024), which employs an annealed guidance strategy to improve trajectory adherence.
-Table 1: Quantitative results on static 3D and dynamic 4D benchmarks.
-We evaluate on public datasets such as LLFF and MipNeRF360, as well as Internet images and videos.
-For static scenes we report $\mathrm{CLIP}_{\mathrm{sim}}$ and FID, and for dynamic scenes $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$ and FVD.
-Metric details are provided in Appendix B.
-Our method consistently surpasses SOTA baselines in generation quality.
-Static Dynamic
-FID $\downarrow$ $\mathrm{CLIP}_{\mathrm{sim}}\uparrow$ FVD $\downarrow$ $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}\uparrow$
+For single-view 3D scene generation, we use data from LLFF (Mildenhall et al., 2019), Tanks and Temples (Knapitsch et al., 2017), MipNeRF 360 (Barron et al., 2022), and diverse internet, real-world, and AI-generated images.
+We report FID (Heusel et al., 2017) and $\mathrm{CLIP}_{\mathrm{sim}}$ (Radford et al., 2021).
+For 4D trajectory control, we use challenging real-world videos, reporting FVD (Unterthiner et al., 2018) and $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$.
+For both tasks, trajectory accuracy is measured with Absolute Trajectory Error (ATE), Relative Pose Error—Translation (RPE-T), and Relative Pose Error—Rotation (RPE-R).
+Full metric definitions are in Appendix B.
+4.2 3D and 4D Trajectory-Controlled Generation
+We compare our method against state-of-the-art baselines on both 3D static scene generation and 4D dynamic video control.
+For 3D novel view synthesis, we evaluate against both training-based (Yu et al., 2024c; 2025; Xiao et al., 2025; Ma et al., 2025a) and training-free (You et al., 2025; Liu et al., 2024) methods.
+For 4D trajectory control, baselines include ReCamMaster (Bai et al., 2025a) and others (Yu et al., 2025; Xiao et al., 2025; Liu et al., 2024).
+Under identical evaluation settings, our training-free method consistently achieves superior results.
+On 3D static scenes, it outperforms both training-based and training-free baselines on public datasets (Fig.
+3, Table 1).
+On 4D clips with challenging camera paths (e.g., arcs, dolly zooms), it yields higher visual fidelity, tighter trajectory alignment, and more coherent scene completion, matching or surpassing costly training-based approaches.
+In both settings, our method plausibly reconstructs unseen regions where baselines often produce distortions.
+Our approach particularly excels in difficult cases.
+It handles human-centric scenes, which require high consistency, and can synthesize photorealistic and structurally consistent $360^{\circ}$ views from a single input (Appendix E.5).
+By preserving model priors, it strikes a strong balance between controllability and fidelity.
+We tested our method on a lighter SVD model (Blattmann et al., 2023) and still achieved high visual quality, confirming its strong performance across different model scales (Appendix E.2).
+Beyond benchmarks, the framework serves as a versatile tool for video post-production.
+It can stabilize videos by smoothing camera motion, control paths for localized super-resolution or outpainting, and perform masked edits like object addition/removal or subject try-on effects.
+These capabilities highlight its utility for real-world video re-rendering (more results in Appendix E.4).
+Table 1: Quantitative comparison with SOTA methods on 3D static and 4D dynamic scenes, using public and internet data.
+We evaluate generation quality (FID, $\mathrm{CLIP}_{\mathrm{sim}}$ for static; FVD, $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$ for dynamic) and trajectory accuracy (ATE, RPE-T, RPE-R).
+All methods use official code with identical inputs.
+$\uparrow$: Higher is better, $\downarrow$: Lower is better.
+Our method achieves the best or second-best results on all metrics.
+Metric details are in Appendix B.
+Generation Quality Trajectory Accuracy
+Static Dynamic Static Dynamic
+FID $\downarrow$ $\mathrm{CLIP}_{\mathrm{sim}}\uparrow$ FVD $\downarrow$ $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}\uparrow$ ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$ ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$
 See3D (Ma et al., 2025a)
-123.26 0.941 - -
+123.26 0.941 0.091 0.089 0.250
 ViewCrafter (Yu et al., 2024c)
-117.50 0.930 - -
+117.50 0.930 0.236 0.315 0.728
 ViewExtrapolator (Liu et al., 2024)
-125.50 0.930 108.48 0.913
+125.50 0.930 108.48 0.913 0.183 0.260 0.882 1.040 1.208 4.750
 TrajectoryAttention (Xiao et al., 2025)
-122.37 0.920 106.94 0.911
+122.37 0.920 106.94 0.911 0.159 0.238 0.532 0.605 1.238 3.560
 TrajectoryCrafter (Yu et al., 2025)
-111.49 0.910 97.31 0.923
+111.49 0.910 97.31 0.923 0.090 0.152 0.267 0.431 1.078 8.950
 NVS-Solver (You et al., 2025)
-118.64 0.937 - -
+118.64 0.937 0.224 0.268 1.056
 WorldForge (Ours)
-96.08 0.948 93.17 0.938
-Table 2: Quantitative comparison of camera-trajectory accuracy on static and dynamic scenes.
-We evaluate against SOTA methods using ATE, RPE-T, and RPE-R, lower is better.
-Across both settings, our method achieves the best or second-best results on all metrics.
-Metric definitions and computation details are provided in Appendix B.
-Static Dynamic
-ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$ ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$
-See3D (Ma et al., 2025a)
-0.091 0.089 0.250 - - -
-ViewCrafter (Yu et al., 2024c)
-0.236 0.315 0.728 - - -
-ViewExtrapolator (Liu et al., 2024)
-0.183 0.260 0.882 1.040 1.208 4.750
-TrajectoryAttention (Xiao et al., 2025)
-0.159 0.238 0.532 0.605 1.238 3.560
-TrajectoryCrafter (Yu et al., 2025)
-0.090 0.152 0.267 0.431 1.078 8.950
-NVS-Solver (You et al., 2025)
-0.224 0.268 1.056 - - -
-WorldForge (Ours)
-0.077 0.086 0.221 0.527 0.826 2.690
-[…]
-Frames Resolution
-Inference Time
-Base Video
-Model
-Training-
-Free
-See3D (Ma et al., 2025a)
-25 $576\times 1024$ 1.7 Custom
-ViewCrafter (Yu et al., 2024c)
-25 $576\times 1024$ 1.8 Custom
-ViewExtrapolator (Liu et al., 2024)
-25 $576\times 1024$ 1.6 SVD
-TrajectoryAttention (Xiao et al., 2025)
-25 $76\times 1024 5.5$ SVD
-TrajectoryCrafter (Yu et al., 2025)
-25 $384\times 672$ 1.7 CogVideoX
-NVS-Solver (You et al., 2025)
-25 $576\times 1024$ 9.3 SVD
-ReCamMaster (Bai et al., 2025a)
-81 $480\times 832$ 14.6 Wan 2.1 T2V
-WorldForge (Ours, 720P)
-25 $1280\times 720$ 17.3 Wan 2.1 I2V
-WorldForge (Ours, 480P)
-25 $480\times 832$ 6.8 Wan 2.1 I2V
-WorldForge (Ours, on SVD)
-25 $576\times 1024$ 1.3 SVD
-We evaluate all methods under consistent settings with the same input image, camera trajectory, depth, and target scenes across LLFF (Mildenhall et al., 2019), MipNeRF-360 (Barron et al., 2022), and Tanks-and-Temples (Knapitsch et al., 2017).
-Results (Figs.3, Table1) show that our inference-time guidance consistently outperforms both training-based and training-free baselines.
-Moreover, we further assess human-centric scenes, which are more sensitive to visual consistency and content quality.
-As shown in Fig.
-4, prior methods often fail in such cases.
-In contrast, our approach, supported by the proposed guidance strategy, preserves model priors and faithfully follows the target trajectory.
-This offers an effective paradigm for reconciling controllability with generalization and high fidelity.
-To independently validate the effect of our guidance strategy, we further evaluate on the lighter U-Net-based SVD model (Blattmann et al., 2023); even in this setting, our method attains superior visual quality.
-Figure 4: Static 3D generation on human-centric scenes.
-Existing methods struggle, particularly with motion-prone shots (left) and portrait close-ups (right).
-On the left, baselines introduce artifacts and unintended motion.
-On the right, most fail to produce plausible results; TrajectoryCrafter (Yu et al., 2025) recovers coarse structure but lacks detail and visual appeal.
-In contrast, our method maintains scene stationarity under trajectory guidance and produces natural, faithful renderings, achieving both precise control and high perceptual quality.
-Notably, our framework can synthesize high-quality, photorealistic, and structurally consistent $360^{\circ}$ views from a single input without relying on panoramic intermediates (Fig.
-5).
-This capability extends to complex outdoor and open-world scenes, pushing the limits of single-view novel view synthesis.
-Figure 5: 360° orbit views from a single real-world outdoor image.
-With precise trajectory control and realistic rendering, our method overcomes the viewpoint limitation of single-image generation and produces ultra-wide views of complex real scenes.
-Unlike panorama-based approaches, it directly supports object-centric trajectories and achieves higher visual quality.
-4.3 4D Trajectory Control
-For dynamic trajectory control, we compare against leading video-to-video models.
-These include ReCamMaster (Bai et al., 2025a), which employs a sophisticated conditioning mechanism trained on large-scale synthetic data, TrajectoryCrafter (Yu et al., 2025), with its dual-stream diffusion architecture, TrajectoryAttention (Xiao et al., 2025), which processes motion through an additional attention module, and ViewExtrapolator (Liu et al., 2024).
-Following the protocols of prior works, we evaluate on diverse and challenging video clips, including arcs, dolly zooms, and composite paths.
-Our training-free method consistently delivers higher visual fidelity, better trajectory alignment, and more coherent scene completion.
-As shown in Table 1 and Fig.
-6, it matches or surpasses state-of-the-art methods that require costly training, demonstrating the strength of leveraging latent 3D priors in existing models.
-Notably, our approach excels at plausibly reconstructing unseen regions, while other methods often produce distortions or implausible content.
-Figure 6: Comparison of 4D trajectory-controlled re-rendering.
-Baselines often produce implausible artifacts (e.g., flattened faces, floating heads), reflecting limited use of pretrained priors.
-Our inference-time guidance leverages these latent world priors to re-render realistic, high-quality content along the target trajectory.
-We compare against state-of-the-art baselines under identical inputs; for ReCamMaster (text-controlled), parameters are adjusted to match the target path.
-Beyond benchmarks, the framework supports various post-production tasks.
-It can stabilize videos by smoothing camera motion and control camera paths to enable localized super-resolution or outpainting.
-In addition, by specifying masked regions, it supports creative video content edits such as object addition or removal, subject replacement, and try-on effects (see Fig.
-7).
-These capabilities highlight its versatility for real-world video re-rendering.
-Figure 7: Other video effects enabled by our method.
-Beyond video re-cam, our flexible depth-based warping also supports various video editing operations, such as freezing the camera, stabilizing video, and editing video content.
-These extensions further broaden the practical scope of our approach.
-4.4 Ablation Experiments
-We conducted ablation studies to assess the contribution of each component in our framework.
+96.08 0.948 93.17 0.938 0.077 0.086 0.221 0.527 0.826 2.690
 Figure 4: Ablation of the proposed components.
@@
 The full model achieves the best fidelity and control, demonstrating their complementary effects.
+4.3 Ablation Experiments
+We ablate each module and design choice in our framework.
 Component Analysis.
-We examined the effects of removing Intra-Step Recursive Refinement (IRR), Flow-Gated Latent Fusion (FLF), and Dual-Path Self-Corrective Guidance (DSG).
-As shown in Fig.
-8 without IRR, trajectory guidance cannot be injected during inference, and the model fails to follow the target path.
-Removing FLF, i.e., not separating motion and appearance channels, damages the model’s priors and produces unnatural results.
-Without DSG, noise from warped trajectories propagates into the generation, causing low quality and artifacts.
-The full model achieves the best results, confirming that these components work synergistically to ensure robust and precise control.
-Base Model.
-We replaced the Wan2.1 (Wan et al., 2025) model with the U-Net-based SVD (Blattmann et al., 2023) model to verify model-agnosticism and transferability (See in Fig.
-9).
-Fair comparisons with other SVD-based methods show that our approach integrates seamlessly into pre-trained SVD.
-Although limited by SVD’s weaker priors, our method still reaches state-of-the-art performance within the same model.
-This demonstrates that our guidance is architecture-independent and suggests even greater potential with stronger future base models.
-Depth Estimation Model.
-As shown in Fig.
-10 We tested VGGT (Wang et al., 2025b), Mega-SaM (Li et al., 2025), UniDepth (Piccinelli et al., 2024) and DepthCrafter (Hu et al., 2025) for depth-based warping.
-Each shows distinct strengths and weaknesses, yet our method remains effective across all.
-Thanks to the generative model’s strong 3D prior, many warping artifacts (e.g., tearing, stretching) are automatically corrected during refinement.
-This robustness enables plug-and-play integration with various SOTA depth estimation techniques without sacrificing performance.
-Figure 9: Ablation across different VDMs.
-To rule out the influence of the intrinsic performance advantage of the VDM (Wan2.1 (Wan et al., 2025)) and to verify the method’s transferability, we port the proposed guidance to a compact U-Net–based SVD model (Blattmann et al., 2023) and compare against SVD-based SOTA baselines.
-Experiments show that the guidance transfers seamlessly, makes the native SVD controllable, and achieves SOTA performance in content quality, structural plausibility, and trajectory consistency.
-Figure 10: Depth estimation models ablation.
-Our method leverages the inherent world knowledge of VDMs to correct errors and fill missing regions even under challenging inputs (left).
-This strong self-correction ability ensures broad compatibility with different depth estimators (right).
-Despite variations or noise in depth-based warping, it reliably compensates through learned priors and produces realistic, high-quality results.
-5 Conclusion and Limitation
-We present WorldForge, a training-free framework for trajectory-controllable generation in both static 3D scenes and dynamic 4D scenes.
-Our method tackles the persistent challenge in controllable video generation of balancing high visual quality, strong generalization, and precise controllability.
-At its core is a unified guidance strategy—Intra-Step Recursive Refinement (IRR), Flow-Gated Latent Fusion (FLF), and Dual-Path Self-Corrective Guidance (DSG)—that operates entirely at inference time.
-By decoupling motion from appearance features and correcting trajectory drift caused by structural noise, it injects fine-grained predefined trajectory constraints while preserving the valuable prior world knowledge embedded in the base model.
-Extensive experiments demonstrate state-of-the-art performance on both 3D and 4D trajectory generation tasks, offering a new paradigm for exploring spatial intelligence and emergent world models in large-scale generative systems.
-While our framework can repair many distortions introduced by depth-based warping, it may fail under extremely poor depth estimations (e.g., completely flattened subjects or severe foreground–background entanglement).
-Moreover, due to the global nature of our guidance, control over small objects or fine details remains limited.
-Future work will explore integrating fine-grained control mechanisms and applying our approach to more powerful generative models.
+We remove IRR, FLF, and DSG in turn (Fig.
+4).
+Removing IRR disables trajectory guidance at inference time, resulting in failure to follow the target path.
+Without FLF, i.e., lacking motion/appearance separation, model priors become entangled, leading to unnatural outputs.
+Removing DSG introduces noise from warped trajectories into the generation process, causing artifacts and degrading visual quality.
+The complete model yields the best results, showing that all components are essential and work synergistically to enable robust and precise control.
+Video Model and Depth Model.
+We replace Wan 2.1 (Wan et al., 2025) with the U-Net–based SVD (Blattmann et al., 2023) to test model-agnosticism.
+It consistently achieved strong results, demonstrating its effectiveness across a range of model architectures(Appendix E.2).
+Similarly, our method remains robust regardless of the warping model.
+We experimented with different depth estimators (Wang et al., 2025b; Li et al., 2025; Piccinelli et al., 2024; Hu et al., 2025) and found that the VDM’s strong 3D priors effectively mitigate many warping artifacts.
+This allows for a plug-and-play integration with various depth models (Appendix E.3).
+5 Conclusion
+We present WorldForge, a training-free framework for trajectory-controllable generation in static 3D and dynamic 4D scenes.
+Our method effectively balances visual quality, generalization, and precise control in video synthesis.
+At its heart is a unified, inference-time guidance strategy—comprising Intra-Step Recursive Refinement (IRR), Flow-Gated Latent Fusion (FLF), and Dual-Path Self-Corrective Guidance (DSG).
+By decoupling motion from appearance and correcting trajectory drift, our framework injects fine-grained control while preserving the rich world priors of the base model.
+Extensive experiments show state-of-the-art performance on both 3D and 4D generation tasks, offering a new path for exploring spatial intelligence in large-scale generative systems.
+While our framework corrects many warping-induced distortions, it can fail with severely inaccurate depth estimations (e.g., completely flattened subjects or severe foreground–background entanglement).
+Furthermore, the global nature of our guidance offers limited control over small objects and fine details.
+Future work will focus on integrating fine-grained control mechanisms and applying our method to more powerful generative models.
 Appendix A Proof of the equivalence between diffusion and flow models
@@
 FID (Heusel et al., 2017) measures image generation quality by comparing the distribution of real and generated images in the Inception-V3 feature space.
-We use an ImageNet-pretrained Inception-V3 model and extract 2048-dimensional features from the pool3 layer.
+We use an ImageNet-pretrained Inception-V3 (Szegedy et al., 2016) model and extract 2048-dimensional features from the pool3 layer.
 The FID score is computed as:
@@
 Fréchet Video Distance (FVD).
-FVD (Unterthiner et al., 2018) extends FID to the video domain by measuring distribution differences in spatio-temporal feature space.
-For computational efficiency, we implement a simplified version that treats video frame sequences as image collections and applies Inception-V3 frame-wise feature extraction.
-The FVD score follows the same Fréchet distance formula as FID but operates on video frame features.
+FVD (Unterthiner et al., 2018) measures distributional differences between real and generated video using pretrained spatio-temporal features.
+We use an I3D (Inflated 3D ConvNet) pretrained on Kinetics (Carreira & Zisserman, 2017) and extract 1024-D features from the global average pooling layer for each video clip.
+Following FID, we compute the Fréchet distance between the Gaussian fits of real and generated I3D features:
+$\text{FVD}=\|\mu_{r}-\mu_{g}\|_{2}^{2}+\mathrm{Tr}\!\left(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{1/2}\right),$ (21)
+with $\mu_{r},\mu_{g}$ and $\Sigma_{r},\Sigma_{g}$ estimated over clip-level I3D features.
 Video CLIP Similarity ($\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$).
@@
 Absolute Trajectory Error (ATE).
-ATE measures the global consistency between estimated and reference camera trajectories by computing Euclidean distances between corresponding camera positions.
-The ATE for each timestamp is computed as:
-$\text{ATE}_{i}=\|\mathbf{t}_{\text{ref},i}-\mathbf{t}_{\text{est},i}\|_{2}$ (28)
-where $\mathbf{t}_{\text{ref},i}$ and $\mathbf{t}_{\text{est},i}$ are the 3D position vectors of the reference and estimated trajectories at timestamp $i$, respectively.
-The root mean square error over all $n$ trajectory points is:
-$\text{ATE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\text{ATE}_{i}^{2}}$ (29)
-Relative Pose Error - Translation (RPE-T).
-RPE-T evaluates the local accuracy of camera translation between consecutive frames, reflecting short-term trajectory consistency.
-For consecutive frames, the relative translation error is:
-$\text{RPE-T}_{i}=\|\Delta\mathbf{t}_{\text{ref},i}-\Delta\mathbf{t}_{\text{est},i}\|_{2}$ (30)
-where $\Delta\mathbf{t}_{\text{ref},i}=\mathbf{t}_{\text{ref},i+1}-\mathbf{t}_{\text{ref},i}$ and $\Delta\mathbf{t}_{\text{est},i}=\mathbf{t}_{\text{est},i+1}-\mathbf{t}_{\text{est},i}$.
-The RMSE over all n1n-1 consecutive frame pairs is:
-$\text{RPE-T}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n-1}\text{RPE-T}_{i}^{2}}$ (31)
-Relative Pose Error - Rotation (RPE-R).
-RPE-R assesses the accuracy of camera orientation changes between consecutive frames.
-For rotation matrices $\mathbf{R}_{\text{ref},i}$ and $\mathbf{R}_{\text{est},i}$, the angular difference in degrees is computed as:
-$\text{RPE-R}_{i}=\arccos\left(\frac{\text{trace}(\Delta\mathbf{R}_{\text{ref},i}^{T}\Delta\mathbf{R}_{\text{est},i})-1}{2}\right)\cdot\frac{180}{\pi}$ (32)
-where $\Delta\mathbf{R}_{\text{ref},i}=\mathbf{R}_{\text{ref},i+1}\mathbf{R}_{\text{ref},i}^{T}$ and $\Delta\mathbf{R}_{\text{est},i}=\mathbf{R}_{\text{est},i+1}\mathbf{R}_{\text{est},i}^{T}$.
-The RMSE over all consecutive frame pairs is:
-$\text{RPE-R}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n-1}\text{RPE-R}_{i}^{2}}$ (33)
+Before evaluation, we align the estimated trajectory to the reference by a global Sim3 transform (scale, rotation, translation).
+Let the aligned pose components be $\tilde{\mathbf{t}}_{\text{est},i}$ and $\tilde{\mathbf{R}}_{\text{est},i}$.
+ATE measures global consistency by the Euclidean distance between corresponding camera positions:
+$\text{ATE}_{i}=\big\|\mathbf{t}_{\text{ref},i}-\tilde{\mathbf{t}}_{\text{est},i}\big\|_{2},\qquad\text{ATE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\text{ATE}_{i}^{2}}.$ (23)
+Relative Pose Error — Translation (RPE-T). RPE-T evaluates local translation accuracy between consecutive frames. Define relative motions via poses (index gap $\Delta\!=\!1$):
+$\Delta\mathbf{T}_{\text{ref},i}=\mathbf{T}_{\text{ref},i}^{-1}\mathbf{T}_{\text{ref},i+1},\qquad\Delta\mathbf{T}_{\text{est},i}=\tilde{\mathbf{T}}_{\text{est},i}^{-1}\tilde{\mathbf{T}}_{\text{est},i+1}.$ (24)
+Let $\Delta\mathbf{t}_{\text{ref},i}$ and $\Delta\mathbf{t}_{\text{est},i}$ be the translation parts of these relative transforms. The per-step error and RMSE are:
+$\text{RPE-T}_{i}=\big\|\Delta\mathbf{t}_{\text{ref},i}-\Delta\mathbf{t}_{\text{est},i}\big\|_{2},\qquad\text{RPE-T}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n-1}\text{RPE-T}_{i}^{2}}.$ (25)
+Relative Pose Error — Rotation (RPE-R). RPE-R assesses the accuracy of orientation changes between consecutive frames. Let the relative rotations be
+$\Delta\mathbf{R}_{\text{ref},i}=\mathbf{R}_{\text{ref},i}^{-1}\mathbf{R}_{\text{ref},i+1},\qquad\Delta\mathbf{R}_{\text{est},i}=\tilde{\mathbf{R}}_{\text{est},i}^{-1}\tilde{\mathbf{R}}_{\text{est},i+1}.$ (26)
+The per-step angular error (degrees) and RMSE are:
+$\text{RPE-R}_{i}=\arccos\!\left(\frac{\mathrm{trace}\!\big(\Delta\mathbf{R}_{\text{ref},i}^{\top}\Delta\mathbf{R}_{\text{est},i}\big)-1}{2}\right)\cdot\frac{180}{\pi},\qquad\text{RPE-R}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n-1}\text{RPE-R}_{i}^{2}}.$ (27)
 B.4 Implementation Details
 Preprocessing.
-Images are resized to $299\times 299$ pixels for FID/FVD computation and $224\times 224$ pixels for CLIP-based metrics, with appropriate normalization applied.
-For camera trajectory evaluation, images are resized to $720\times 480$ pixels and uniformly sampled to 20 frames while preserving the first and last frames.
-Videos are uniformly sampled to 25 frames while preserving the first and last frames.
-For computational efficiency, FVD calculation further samples to 16 frames, and $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$ processing is limited to 20 frames for long videos.
+For FID, images are resized to $299\times 299$ and fed to Inception-V3 with standard ImageNet normalization.
+For FVD and CLIP-based metrics, frames are resized to $224\times 224$ with the respective model normalizations.
+For camera trajectory evaluation, images are resized to $720\times 480$ and uniformly sampled to $20$ frames while preserving the first and last frames.
+Videos are uniformly sampled to $25$ frames with first/last preserved; for FVD, we further sample to $16$ frames per clip; for $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$ on long videos, we cap processing at $20$ frames.
 Evaluation Protocol.
-For static scenes with multiple reference images, we directly construct the real distribution using all available images.
-For single-image scenes, we apply minimal augmentation strategies to avoid singular covariance matrices.
-Dynamic scenes maintain frame correspondence between generated and reference videos to ensure fair comparison.
-For camera trajectory evaluation, we employ Structure-from-Motion (SfM) to reconstruct camera poses from image sequences, then apply Sim3 alignment to handle scale ambiguity inherent in monocular reconstruction.
-Trajectory comparisons are performed using the evo toolkit with appropriate alignment and scale correction parameters.
+For static scenes with multiple references, we estimate the real distribution from all reference images; for single-image scenes, we apply minimal augmentation to avoid singular covariances.
+Dynamic scenes maintain frame correspondence between generated and reference videos.
+For trajectories, poses are recovered by SfM, the estimated trajectory is aligned to the reference by Sim3 to resolve scale, and metrics are computed using evo with alignment and scale correction enabled.
+Appendix C Details for FLF Scoring and Settings
+This section provides a detailed breakdown of the Flow-Gated Latent Fusion (FLF) module, which is introduced in Section 3.3.
+The goal of FLF is to identify and selectively update latent channels that are highly relevant to motion, thereby preserving visual details encoded in appearance-focused channels.
+To achieve this, at each denoising step $i$, FLF computes a motion similarity score $S^{(t,c)}$ for each latent channel $c$.
+Below, we detail how this score is calculated.
+Optical Flow Computation
+At each denoising step $i$, we compute optical flow maps for each channel $c$ of both the predicted latent $\hat{\mathbf{x}}_{0}^{(t)}$ and the target trajectory latent $\mathbf{Z}_{\text{traj}}$.
+The computation is performed frame-by-frame; that is, for each latent tensor, we calculate the dense optical flow between consecutive temporal frames using the Farnebäck algorithm (Farnebäck, 2003).
+This process yields a predicted flow map, $\mathcal{F}^{(t,c)}_{\text{pred}}$, and a ground-truth (GT) flow map, $\mathcal{F}^{(t,c)}_{\text{gt}}$.
+At each pixel, the flow is a 2D vector $(u_{*},v_{*})$ representing horizontal and vertical displacement.
+All subsequent metric calculations are performed over the set of valid (i.e., non-occluded) pixels, defined as $\Omega^{(t,c)}=\{(x,y,\tau)\mid\mathbf{M}^{(c)}(x,y,\tau)=1\}$, where $(x,y)$ are pixel coordinates and $\tau$ is the frame index.
+Since optical flow is computed between adjacent frames, for a latent tensor with $T_{l}$ total frames, the index $\tau$ ranges from $1$ to $T_{l}-1$.
+Metric Calculation
+The motion score $S^{(t,c)}$ is derived from three standard optical flow metrics that quantify the error between the predicted flow $\mathcal{F}^{(t,c)}_{\text{pred}}$ and the ground-truth flow $\mathcal{F}^{(t,c)}_{\text{gt}}$ at each step $i$.
+  • Masked End-point Error (M-EPE) measures the average Euclidean distance between the predicted and GT flow vectors over all valid pixels:
+    $\text{M-EPE}^{(t,c)}=\frac{1}{|\Omega^{(t,c)}|}\sum_{(x,y,\tau)\in\Omega^{(t,c)}}\left\|\mathcal{F}^{(t,c)}_{\text{pred}}(x,y,\tau)-\mathcal{F}^{(t,c)}_{\text{gt}}(x,y,\tau)\right\|_{2}.$ (28)
+  • Masked Angular Error (M-AE) calculates the average angular difference in radians between the flow vectors:
+    $\text{M-AE}^{(t,c)}=\frac{1}{|\Omega^{(t,c)}|}\sum_{(x,y,\tau)\in\Omega^{(t,c)}}\arccos\left(\frac{\mathcal{F}^{(t,c)}_{\text{pred}}(x,y,\tau)\cdot\mathcal{F}^{(t,c)}_{\text{gt}}(x,y,\tau)}{\|\mathcal{F}^{(t,c)}_{\text{pred}}(x,y,\tau)\|\cdot\|\mathcal{F}^{(t,c)}_{\text{gt}}(x,y,\tau)\|}\right).$ (29)
+  • Outlier Percentage (Fl-all) is the percentage of pixels in $\Omega^{(t,c)}$ where the flow estimation is considered erroneous.
+Following standard benchmarks, a pixel is flagged as an outlier if its M-EPE exceeds 3 pixels or if its relative error, $\|\mathcal{F}^{(t,c)}_{\text{pred}}-\mathcal{F}^{(t,c)}_{\text{gt}}\|_{2}/\|\mathcal{F}^{(t,c)}_{\text{gt}}\|_{2}$, is greater than 5%.
+We denote this outlier percentage as $\text{F}^{(t,c)}$.
+Normalization and Weighting
+The three metrics exist on different scales, so we first normalize each to the range $[0,1]$ before combining them. This corresponds to the ${\text{Norm}_{k}}^{(t,c)}$ terms used in the main text:
+${\text{Norm}_{\text{E}}}^{(t,c)}$ $=\min(\text{M-EPE}^{(t,c)}/n_{\text{E}},1),$ (30)
+${\text{Norm}_{\text{A}}}^{(t,c)}$ $=\min(\text{M-AE}^{(t,c)}/n_{\text{A}},1),$
+${\text{Norm}_{\text{F}}}^{(t,c)}$ $=\min(\text{F}^{(t,c)}/n_{\text{F}},1),$
+where $n_{\text{E}},n_{\text{A}},$ and $n_{\text{F}}$ are normalization constants chosen to reflect typical value ranges for each metric.
+The final motion score $S^{(t,c)}$ is a weighted sum of the inverted normalized errors, as defined in Eq.
+(6) in the main text.
+The weights $\gamma_{k}$ (where $k\in\{\text{E},\text{A},\text{F}\}$ and $\sum_{k}\gamma_{k}=1$) and the normalization constants are set based on common practices in optical flow evaluation to balance each metric’s contribution.
+In our experiments, we set $n_{\text{E}}=10$, $n_{\text{A}}=30$, and $n_{\text{F}}=0.5$.
+The weights in Eq.
+(6) are set to $(\gamma_{\text{E}},\gamma_{\text{A}},\gamma_{\text{F}})=(0.4,0.3,0.3)$.
+Appendix D Visual Comparison of DSG and Naive CFG
+In contrast, our proposed DSG mechanism successfully balances trajectory control and visual fidelity.
+The images generated with DSG maintain structural integrity and high perceptual quality while closely following the intended camera path.
+This comparison validates our central claim that a specialized guidance mechanism like DSG is necessary for robust control in our warp-and-repaint framework and demonstrates its clear superiority over a direct application of CFG.
+As discussed in the main text, our Dual-Path Self-Corrective Guidance (DSG) is specifically designed to handle the large angular difference between the guided velocity ($\mathbf{v}_{t}^{\text{traj}}$) and the unguided velocity ($\mathbf{v}_{t}^{\text{ori}}$).
+This large divergence, with observed cosine similarities often between 0.3–0.6, renders a naive application of the standard Classifier-Free Guidance (CFG) formula ineffective, as CFG is designed for scenarios where the conditional and unconditional paths are closely aligned.
+To visually demonstrate this, we conduct a direct comparison, shown in Fig.
+5.
+We compare our full DSG framework against both a standard CFG implementation and an ablated version of our method without the adaptive weight $\beta_{t}$.
+As the results show, applying a naive CFG formulation leads to severe visual artifacts and distorted object structures, while removing the adaptive weight also reduces guidance stability.
+Figure 5: Visual comparison of our proposed DSG and naive CFG for trajectory guidance.In many test scenes, naive CFG fails to handle the large angular difference between the two velocity fields, resulting in significant artifacts and errors.
+Removing the adaptive weighting factor $\beta_{t}$ from our method (DSG w/o $\beta_{t}$) leads to reduced guidance stability and introduces errors.
+In contrast, our full DSG framework stably generates high-fidelity and structurally consistent results.
+Appendix E More Experimental Results
+This section provides additional experiments and results that complement the findings presented in the main paper.
+We include a detailed efficiency analysis, further ablation studies, and more qualitative examples to fully demonstrate the capabilities and robustness of our framework.
+E.1 Efficiency and Runtime Analysis
+As mentioned in Section 4.1, our framework is training-free and operates entirely at inference time.
+Table 2 provides a detailed comparison of inference efficiency against several state-of-the-art methods on a single NVIDIA A100 GPU.
+Our method incurs zero training cost, which is a significant advantage over approaches that require costly fine-tuning.
+The primary computational overhead comes from the Intra-Step Recursive Refinement (IRR) module, which increases the inference time by approximately 40–50% compared to running only the base VDM.
+Despite this, our framework achieves inference speeds that are comparable to, and in some cases faster than, prior methods.
+For instance, when integrated with the SVD backbone, our method is more efficient than several other SVD-based baselines.
+This analysis demonstrates that our framework achieves strong controllability without prohibitive computational costs, offering an efficient and effective alternative to training-intensive pipelines.
+[…]
+Frames Resolution
+Inference Time (min)
+Base Video Model
+Training-Free
+See3D (Ma et al., 2025a)
+25 $576\times 1024$ 1.7 Custom
+ViewCrafter (Yu et al., 2024c)
+25 $576\times 1024$ 1.8 Custom
+ViewExtrapolator (Liu et al., 2024)
+25 $576\times 1024$ 1.6 SVD
+TrajectoryAttention (Xiao et al., 2025)
+25 $76\times 1024 5.5$ SVD
+TrajectoryCrafter (Yu et al., 2025)
+25 $384\times 672$ 1.7 CogVideoX
+NVS-Solver (You et al., 2025)
+25 $576\times 1024$ 9.3 SVD
+ReCamMaster (Bai et al., 2025a)
+81 $480\times 832$ 14.6 Wan 2.1 T2V
+WorldForge (Ours, 720P)
+25 $720\times 1280$ 17.3 Wan 2.1 I2V
+WorldForge (Ours, 480P)
+25 $480\times 832$ 6.8 Wan 2.1 I2V
+WorldForge (Ours, on SVD)
+25 $576\times 1024$ 1.3 SVD
+E.2 Ablation on Video Diffusion Models
+To verify that our performance is due to our proposed guidance mechanism and not just the power of the primary VDM (Wan2.1), we conducted an ablation study by porting our entire framework to the widely-used, U-Net-based Stable Video Diffusion (SVD) model (Blattmann et al., 2023).
+We made a minor adjustment to adapt to the characteristics of SVD and its native EDM sample.
+We then performed a fair comparison against other state-of-the-art methods that are also built on the SVD backbone, using identical inputs.
+The results, shown in Fig.
+6, demonstrate that our guidance transfers seamlessly to this new architecture.
+It makes the native SVD model controllable and capable of following specific trajectories, achieving SOTA performance among SVD-based methods in content quality, structural plausibility, and trajectory consistency.
+This experiment confirms that our guidance is architecture-independent and suggests it will have strong potential when paired with future, more powerful base models.
+Figure 6: Ablation across different VDMs.
+To rule out the influence of the intrinsic performance advantage of the VDM (Wan2.1 (Wan et al., 2025)) and to verify the method’s transferability, we port the proposed guidance to a compact U-Net–based SVD model (Blattmann et al., 2023) and compare against SVD-based SOTA baselines.
+Experiments show that the guidance transfers seamlessly, makes the native SVD controllable, and achieves SOTA performance in content quality, structural plausibility, and trajectory consistency.
+E.3 Ablation on Depth Estimation Models
+Our framework relies on a warp-and-repaint strategy, where the quality of the initial warp is dependent on a depth estimation model.
+To test the robustness and flexibility of our approach, we evaluated its performance with several different state-of-the-art depth estimators: VGGT (Wang et al., 2025b), UniDepth (Piccinelli et al., 2024), Mega-SaM (Li et al., 2025), and DepthCrafter (Hu et al., 2025).
+As shown in Fig.
+7, our method demonstrates broad compatibility and maintains high performance across all tested depth models.
+Even when the depth-based warping produces challenging inputs with noise, errors, or significant missing regions (disocclusions), our framework effectively compensates.
+This resilience stems from the strong generative world priors of the underlying VDM, which are leveraged by our guidance modules to correct artifacts and plausibly fill in missing areas during the repainting stage.
+This self-correction capability confirms that our framework can be used in a plug-and-play manner with various depth estimation techniques without sacrificing the quality of the final output.
+Figure 7: Depth-models ablation.
+Our method leverages the inherent world knowledge of VDMs to correct errors and fill missing regions even under challenging inputs (left).
+This strong self-correction ability ensures broad compatibility with different depth estimators (right).
+Despite variations or noise in depth-based warping, it reliably compensates through learned priors and produces realistic, high-quality results.
+E.4 Applications in Video Editing
+Beyond trajectory-controlled generation, our framework’s flexibility makes it a powerful tool for various video post-production and editing tasks.
+This includes effects like video stabilization, camera freezing, and dynamic viewpoint switching.
+Furthermore, by incorporating a flexible masking strategy, our framework can perform diverse content edits such as object removal, addition, subject replacement, and virtual try-on seamlessly.
+The general process for these edits involves first segmenting the target region in each frame using a tool like SAM (Kirillov et al., 2023).
+The desired edit is then applied to the first frame (e.g., using Gemini (Comanici et al., 2025)).
+Finally, this edited frame and the corresponding masks are processed by our pipeline to render a temporally consistent result.
+For adding new objects where none exist in the source video, a simple bounding box can be provided to guide the placement.
+Fig.
+8 shows several qualitative examples of these video editing effects.
+Figure 8: Other video effects enabled by our method.
+Beyond video re-cam, our flexible depth-based warping also supports various video editing operations, such as freezing the camera, stabilizing video, and editing video content.
+These extensions further broaden the practical scope of our approach.
+E.5 Generation on Challenging Scenes
+Our approach demonstrates robust performance in difficult cases where other methods may falter. We highlight two such scenarios: human-centric scenes and single-image 360° view generation.
+Human-Centric Scenes
+Human-centric scenes are challenging for novel view synthesis due to the need for high structural and temporal consistency.
+As shown in Fig.
+9, some methods can struggle with these cases, sometimes introducing artifacts, unintended motion, or difficulty rendering plausible facial features.
+For instance, TrajectoryCrafter (Yu et al., 2025) may recover the coarse structure, but can introduce unnatural facial deformations.
+In contrast, our method’s use of strong generative priors and precise trajectory guidance helps maintain scene stationarity and consistency, producing more natural renderings that better preserve the subject’s appearance.
+Figure 9: Static 3D generation on human-centric scenes.
+Existing methods struggle, particularly with motion-prone shots (left) and portrait close-ups (right).
+On the left, baselines introduce artifacts and unintended motion.
+On the right, most fail to produce plausible results; TrajectoryCrafter (Yu et al., 2025) recovers coarse structure but lacks detail and visual appeal.
+In contrast, our method maintains scene stationarity under trajectory guidance and produces natural, faithful renderings, achieving both precise control and high perceptual quality.
+360° View Generation
+Generating full 360° views from a single image is another demanding task.
+Our framework’s precise trajectory control helps address the limited field-of-view of the source image, enabling the creation of coherent, object-centric orbit views of complex scenes, as shown in Fig.
+10.
+We achieve the full 360° loop by generating a sequence where the final frame seamlessly connects to the first.
+This is made possible by our precise guidance, which maintains high image quality and prevents the accumulation of errors over the entire long-range trajectory, a common point of failure for other methods.
+Unlike traditional panoramic approaches, our method directly generates a continuous view along a given trajectory.
+This can offer more flexibility and strong visual quality, particularly for object-centric paths.
+Figure 10: 360° orbit views from a single real-world outdoor image.
+With precise trajectory control and realistic rendering, our method overcomes the viewpoint limitation of single-image generation and produces ultra-wide views of complex real scenes.
+Unlike panorama-based approaches, it directly supports object-centric trajectories and achieves higher visual quality.